### v1

# STABSim: A Parallelized Clifford Simulator with Features Beyond Direct Simulation

Abstract.
Efficient stabilizer-based computation is a key task in quantum computing as it provides a classically tractable insight into quantum algorithms, which otherwise quickly become impossible at scale on a classical device. Formulations in quantum error correction, hardware verification, hybrid computing, and more, all benefit from the ability of Clifford gates to be accurately evaluated by classical computers. However, current stabilizer simulators have significant limitations in scaling and use-case beyond direct simulation. To address this, we present a GPU-accelerated tableau stabilizer simulator that scales efficiently in direct circuit simulation, and supports a range of stabilizer-based tasks. We show how the simulator is used to quickly calculate Pauli commutation groupings between Pauli strings to reduce the number of circuit runs in quantum chemistry problems. We also present a Clifford+T circuit transpiler based on STABSim, which uses the simulator framework to track relations in non-Clifford rotations as Clifford gates are absorbed into the measurement step of the circuit. The T rotation layers left behind are then further optimized in the simulator structure to create a reduced circuit containing only the minimum non-Clifford rotations, improving the time to simulate and reducing the cost to run on real quantum devices.
copyright: none
1. Introduction
Quantum computing holds the promise of dramatically accelerating solutions to classically hard problems in areas such as cryptography (Shor, 1999; Gisin et al., 2002), optimization (Farhi et al., 2014; Han and Kim, 2002), quantum chemistry (Georgescu et al., 2014; Kandala et al., 2017), and finance (Rebentrost et al., 2018; Woerner and Egger, 2019), etc. In recent years, hardware demonstrations have shown an increasing number of qubits while reducing error rates. Quantum devices capable of demonstrating of quantum advantage in certain tasks have been reported (Arute et al., 2019; Wu et al., 2021; Gao et al., 2025). However, to gain the full potential of quantum computing, the error of quantum systems needs to be carefully managed and corrected (Nielsen and Chuang, 2010) using quantum error correction (QEC) codes.
QEC encodes dozens of physical qubits into one logical qubit and actively detecting and correcting errors that occur in the quantum systems. QEC lays the foundation for fault-tolerant quantum computing. Early demonstrations of small-scale QEC codes have shown encouraging results (Ryan-Anderson et al., 2021; Krinner et al., 2022; Acharya et al., 2023; Bluvstein et al., 2024; Acharya et al., 2025), but the path toward mid-term and far-term devices featuring thousands of logical qubits demands significant advances in architecture, control, and simulation (Gidney and Ekerå, 2021). In particular, designing and verifying different QEC codes such as surface codes, color codes, and especially understanding the error threshold and the noise performance of logical gate operations, requires efficient large-scale classical simulation tools.
Figure 1. Simulation time of surface codes with distances ranging from 3 to 71, using Qiskit method=’stabilizer’ and Stim on an AMD EPYC 7763 processor.
QEC codes primarily consist of Clifford operations, which can be efficiently modeled using the stabilizer formalism (Gottesman, 1998; Nielsen and Chuang, 2010). Unlike general quantum circuit simulators that scale exponentially with qubit count, stabilizer simulators scale polynomially, enabling large-scale QEC design verification. Despite this efficiency, existing stabilizer simulators, including Qiskit (Qiskit Development Team, 2021) and Stim (Gidney, 2021). Qiskit, in addition to its widely used quantum compilation and emulation features, offers a stabilizer simulation backend that can simulate noisy Clifford circuits if all errors in the noise model are also Clifford errors. This has proven beneficial for debugging near-term algorithms and verifying small to mid-sized QEC codes (Goto, 2024; Bravyi et al., 2024a; Cross et al., 2024). Meanwhile, Stim stands out for its meticulous single-threaded CPU optimizations, significantly accelerating large stabilizer circuits through tight memory layouts and specialized bitwise operations (Gidney, 2021). Both Qiskit and Stim highlight the value of specialized stabilizer simulation in quantum computing research, especially for QEC protocols.
However, as we move forward from the Early fault-tolerant era to Megaquop machines (Preskill, 2018; Katabarwa et al., 2024; Preskill, 2025), we expect to expand the simulation probabilities of stabilizer simulators, targeting large-scale multi-logical-qubit code block simulations. However, Stim’s performance often becomes limited by CPU-centric design constraints such as memory bandwidth and the relatively low degree of parallelism available in general-purpose processors. As shown in Figure 1, while Stim performs significantly better than Qiskit, both Qiskit and Stim scale poorly to larger QEC codes.
Large surface-code circuits with repeated rounds of syndrome extraction, fast feedback, and intricate multi-qubit operations—like those in lattice surgery—push the limits of CPU-only approaches. In particular, the sheer volume of stabilizer updates, measurements, and state transformations can quickly overwhelm the available computing resources, resulting in long simulation times and restricting the scale of feasible experiments. In addition, as Stim is optimized towards QEC applications, other applications relying on Clifford and Pauli operations can have difficulties being mapped onto the framework to leverage the power of Stim.
To tackle these challenges, we propose a GPU-accelerated stabilizer simulator that exploits the massive parallelism and high-bandwidth memory of modern Graphics Processing Units (GPUs). Originally designed for graphics workloads, GPUs today provide thousands of concurrent threads and efficient memory subsystems, making them interesting for the large matrix-like operations at the core of QEC simulation. Previous efforts on GPU implementation did not yield performance improvement (Gidney, 2021), citing the low arithmetic intensity of XOR operations at the core of stabilizer simulation. CPUs certainty outperform GPUs for small and sequential arithmetic by design. However, dedicated GPUs have their own bitwise and synchronization capabilities that can safely enable many operations at once, even in the time-order sensitive context of quantum circuit simulation. Furthermore, comparisons and longer arithmetic operations appear frequently during measurement operations, which are the bottleneck of QEC simulation times.
Our approach encodes stabilizer states directly on the GPU, orchestrating operations so that they can be applied in parallel, thereby significantly reducing simulation time and time scaling for larger circuits. This design enables studies of mid-term and far-term QEC protocols, including high-distance surface codes and complex lattice surgery which scales impractically on lightly-threaded CPUs. We still provide lightweight CPU version which performs within the realm of other modern stabilizer simulators. Both versions are flexible with fast stabilizer and Pauli-based computing features beyond direct simulation. As we will show, efficient stabilizer simulation has useful applications for fields beyond QEC, including chemistry simulation and circuit optimization.
This work will detail the following contributions:
  1. (1)
    The first quantum stabilizer simulator to show performance gains via GPU in Clifford+Measurement circuits.
  2. (2)
    Improved scaling among stabilizer simulators as qubit size increases, lowering the time to evaluate a full surface code simulation and measurement after $\sim 2000$ qubits.
  3. (3)
    Integrated Pauli-Based Computing (PBC) capabilities and a direct quantum chemistry use-case shown for VQE.
  4. (4)
    A novel Clifford+T to PBC transpiler using the stabilizer simulator, showing how stabilizer simulation can be leveraged to improve non-Clifford circuits by reducing T-gate state count by 35-70%. This transpiler is the first to our knowledge to automatically convert any static Clifford+T circuit to a PBC circuit using bitwise methods, and the first to perform T-state optimization.
2. Background
2.1. Clifford Gates
At the core of most quantum computing circuits are Pauli operators, which consist of tensor products of Pauli operators $X$, $Y$, $Z$, and identity operator $I$. With the combination of phase factors {±$\{\pm 1,\pm i\}$, Pauli operators form Pauli groups across $n$ qubits.
The Clifford group comprises all unitary operators that map Pauli operators to other Pauli operators under conjugation. In practice, this group is generated by a small set of gates, e.g., single-qubit Hadamard gates $H$, phase gate $S$, and $CX$ gates, represented in computational basis as
(1) $H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix},\,S=\begin{pmatrix}1&0\\ 0&i\end{pmatrix},\,\mathrm{CX}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$
Proved by the Gottesman-Knill theorem, quantum circuits restricted to Clifford gates and Pauli measurements can be simulated on a classical computer in polynomial time with respect to the circuit size and number of qubits (Aaronson and Gottesman, 2004; Gottesman, 1998; Nielsen and Chuang, 2010). Luckily for us, QEC and many other applications are formulated in a way that Clifford gates are sufficient.
2.2. The Stabilizer formalism
A stabilizer of a quantum state $\lvert\psi\rangle$ is an operator $\hat{A}$ such that
(2) $\hat{A}\ket{\psi}=+1\ket{\psi}$
Although stabilizers can be chosen from broader operator sets, Pauli stabilizers (i.e., tensor products of the single-qubit operators I, X, Y, Z) are especially popular due to their simple algebraic structure. A stabilizer group is a commuting (Abelian) subgroup of the $n$-qubit Pauli group whose elements all square to the identity and do not include the negative identity operator $-I$. A state stabilized by a stabilizer group is a stabilizer state. Stabilizer states can be equivalently described by the stabilizers in the stabilizer group, especially the group generators.
The effect of Clifford group operations on stabilizer states can be simulated by tracking the transformation of the corresponding stabilizer group. Pauli measurement can also be tracked using the following rules (Litinski, 2019),
  1. (1)
    if the operator $g$ being measured commutes with the stabilizer group $\mathcal{S}$, the operator $g\in\mathcal{S}$. The state will not change after the measurement, while the measurement outcome depends on the phase of the operator $g$ in the group $G$.
  2. (2)
    if the operator is not in the stabilizer group $\mathcal{S}$, the measurement outcome can be either ±$\pm 1$. The state after the measurement is stabilized by a new group that contains the new measured operator ±$\pm g$ depending on the measurement outcome and all other group generators that commute with $g$ in the original stabilizer group $\mathcal{S}$.
A stabilizer $\mathcal{S}$ represents one possible tensor product of Pauli operators that stabilize a state by equation 2. However, a pure $n$-qubit stabilizer state is stabilized by $n$ commuting and independent Pauli stabilizers. All of these stabilizers can be listed in a table, where rows are the independent stabilizers, and columns are the qubit index.
(3) $\left(\begin{array}[]{c}\sigma_{11}\oplus...\oplus\sigma_{1n}\rightarrow S_{1}\\ \sigma_{21}\oplus...\oplus\sigma_{2n}\rightarrow S_{2}\\ \vdots\\ \sigma_{n1}\oplus...\oplus\sigma_{nn}\rightarrow S_{n}\\ \end{array}\right)$
In order to efficiently track the transformation of Pauli operators, Pauli operators are frequently represented by binary vectors. Each Pauli operator $\sigma$ is represented by two bits, an $x$ portion and a $z$ portion. Pauli $X$ is represented by $(1,0)$, $Z\mapsto(0,1)$, $Y\mapsto(1,1)$, and $I\mapsto(0,0)$. The commutation relation of two Pauli operators represented as $(x_{1},z_{1})$ and $(x_{2},z_{2})$ can also be easily computed using
(4) $(x_{1}\cdot z_{2})\oplus(x_{2}\cdot z_{1}),$
where $\oplus$ is an xor, and a result of 0 means the two Pauli operators commute.
The key purpose of stabilizers in the context of QEC is that errors can easily knock the state out of the stabilizer subspace of the Hilbert space. In other words, our operator $\hat{A}$, or set of operators, will generate a -1 eigenvalue when it should be stabilizing the state, alerting us to an error.
2.3. The CHP Formalism
Aaronson and Gottesman proposed an optimized version of the tableau, specifically for optimizing the measurement operations, known as the CNOT-Hadamard-Phase (CHP) formalism (Aaronson and Gottesman, 2004). For a $n$-qubit state, instead of a single tableau that only tracks the $n$ stabilizer group generators, in CHP formalism, an extra tableau is to added to track the $n$ destabilizers. Specifically, the tableau that are tracked in the CHP formalism can be represented as
(5) $\left(\begin{array}[]{ccc|ccc|c}x_{11}&\cdots&x_{1n}&z_{11}&\cdots&z_{1n}&r_{1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}&z_{n1}&\cdots&z_{nn}&r_{n}\\ \hline\cr x_{(n+1)1}&\cdots&x_{(n+1)n}&z_{(n+1)1}&\cdots&z_{(n+1)n}&r_{n+1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{(2n)1}&\cdots&x_{(2n)n}&z_{(2n)1}&\cdots&z_{(2n)n}&r_{2n}\end{array}\right)$
where each row represents a stabilizer or destabilizer. The first $n$ rows correspond to the stabilizers, while the last $n$ rows are the destabilizers. The columns $r$ represent the sign bits of each Pauli operator. By tracking the destabilizers, it removes the need to do Gaussian elimination on the stabilizer tableau when a Pauli measurement is performed, reducing computational complexity from $O(n^{3})$ to $O(n^{2})$ (Aaronson and Gottesman, 2004). The tableau requires $4n^{2}+n$ bits to store all necessary stabilizers and destabilizers, including a phase bit for each, for CHP operations. Thankfully, this means that a full tableau representation of systems with tens of thousands of qubits still only requires space on the order of MB. For comparison, a state vector or density matrix representation of the state will grow on the order $O(2^{n})$. These simulation methods become infeasible around roughly 50 qubits (Chen et al., 2025).
2.4. Operations on the CHP Tableau
Once the data structure of the tableau is established, we want to perform minimal bitwise operations to simulate the evolution of the tableau under Clifford operations. Clifford gates transform Pauli operators to Pauli operators under conjugation. For example, $HXH^{\dagger}=Z$. From this, some basic bit transforms on the tableau can be calculated that represent the action of Clifford gates on the stabilizers of a system. Most often, all stabilizers in the tableau will receive the same transformation gate-by-gate, meaning Clifford gate operations can be similar to matrix operations. However, the most time consuming operations, measurements, require more logic.
Gate Binary Tableau Operation
Hadamard ($H$) Swap $x$ and $z$; $r=r\oplus(x\cdot z)$
Phase ($S$) $z=z\oplus x$; $r=r\oplus(x\cdot z)$
CNOT ($\mathrm{CX_{c,t}}$) $x_{t}=x_{t}\oplus x_{c}$; $z_{c}=z_{c}\oplus z_{t}$;
$r=r\oplus x_{c}\cdot z_{t}$
Measure ($M$) $\begin{cases}rowsum(i,p)\{\forall i\mid x_{i,q}=1\}&\text{if }random\\ rowsum(temp,i){\forall i>n}&\text{if }deterministic\\ \end{cases}$
Table 1. Clifford gates and their bit operations on the stabilizer tableau. Hadamard, phase and CNOT operations occur on every row of the tableau, at the column that the gate targets. c and t represent the control and target bit of a CNOT operation.
Other Clifford gates are derived from this basic gate set and minimized. Measurement operations in CHP are more complex. Performing measurements is sped up by tracking a corresponding destabilizer tableau, but each measurement still requires two distinct steps that branch depending on if the measurement is random or deterministic.
Algorithm 1 rowsum(h, i) (Aaronson and Gottesman, 2004)
1:function g($x_{1},z_{1},x_{2},z_{2}$)
2:     if $x_{1}=0$ and $z_{1}=0$ then
3:         return 00
4:     else if $x_{1}=1$ and $z_{1}=1$ then
5:         return $z_{2}-x_{2}$
6:     else if $x_{1}=1$ and $z_{1}=0$ then
7:         return $z_{2}\cdot(2x_{2}-1)$
8:     else if $x_{1}=0$ and $z_{1}=1$ then
9:         return $x_{2}\cdot(1-2z_{2})$
10:     end if
11:end function
12:
13:function rowsum($h,i$)
14:     $sum\leftarrow 2r_{h}+2r_{i}$
15:     for $j=1$ to $n$ do
16:         $sum\leftarrow sum+\textsc{g}(x_{ij},z_{ij},x_{hj},z_{hj})$
17:     end for
18:     if $sum\bmod 4=0$ then
19:         $r_{h}\leftarrow 0$
20:     else if $sum\bmod 4=2$ then
21:         $r_{h}\leftarrow 1$
22:     end if
23:     for $j=1$ to $n$ do
24:         $x_{hj}\leftarrow x_{ij}\oplus x_{hj}$
25:         $z_{hj}\leftarrow z_{ij}\oplus z_{hj}$
26:     end for
27:end function
The first step is to search for a stabilizer that does not commute with the measurement. A non-commuting stabilizer indicates a random measurement outcome. In that case, only non-commuting stabilizers are updated. p is the first stabilizer that does not commute with the measurement. rowsum(i,p)$\{i\mid x_{q,i}=1\}$ updates the phase bit of any non-commuting row i using row p. The current stabilizer at p is then stored in the corresponding destabilizer, and the stabilizer at p is projected to the measurement basis, Z in this case. Finally a random phase bit for p, i.e. the measurement outcome, is chosen.
If all stabilizers commute, the measurement outcome needs to be determined. In this case all stabilizers stay the same, but a temporary bit array of size $n+1$ is used to collect all of the stabilizer updates. rowsum(temp,i)$\forall(i>n)$ collects the outcome from every stabilizer in the tableau. Once complete, the phase bit at temp is the measurement outcome.
2.5. Execution of the Stabilizer Simulator
Stabilizer circuits can be split between single qubit Clifford gates, entangling gates, and measurement gates. The time order of all of these gates is critical for correctness. For example, phase and Hadamard gates do not commute, entangling gates need to exchange information at a specific point in time, and measurement gates need all relevant stabilizers to be caught up. Time order is more sensitive in the GPU implementation. If multiple threads read and write back to the same location in data, changes from one thread may not be written back before another thread accesses the data.
In STAB-GPU, every thread is responsible for its own stabilizer row. As such, single qubit gates can be run without synchronization, since there is no risk of another thread updating a given stabilizer before the previous thread finishes an update. Entangling gates rely on and swap information in multiple columns, but since each thread is responsible for dedicated rows, there is no conflict until measurement. Measurement gates require that the circuit execution is caught up before proceeding. Even more costly, execution has to pause while measurement gates perform intermediate calculation that can easily be overwritten by other threads executing later gates. Luckily there is parallelism to be exploited in measurements discussed in section 3.1 to offset the cost of grid-wise synchronization.
Clifford Circuit – GatesSingle Qubit Gate(q)Measurement(q)Entangling Gate(c,t)c, t in the same block?Synchronize BlockSynchronize GridExecute CliffordExecute MeasurementWaitAll gates processed?Updated TableauNoYesM2-Qubit GateYesNo
Figure 2. Gate-by-gate processing of a Clifford circuit into an binary tableau on GPU. Single qubit targets are denoted by q, while c and t denote the control and target qubits of a two-qubit Clifford gate.
2.6. QEC Operations
One of the most important applications using stabilizer simulation is to simulate QEC operations. Specifically, in the syndrome extraction process, the only operations are Clifford gates between syndrome qubits and data qubits and Pauli base measurements on the syndrome qubits. When we inject Pauli errors as a model of actual noise on the physical qubits, the process can be simulated completely in stabilizer formalism.
(a) A distance 3 surface code layout
(b) $X$-checks
(c) $Z$-checks
Figure 3. Planar surface code.
Here, we take plainer surface code as an example. Planar surface code is a topological quantum error-correcting code realized on a two-dimensional grid of physical qubits with open boundary conditions, enabling local stabilizer measurements and scalable fault tolerance. The qubit layout in a distance $5$ surface code patch is shown in Fig. 3(a), where the data qubits (blue) are the qubits encoding the quantum information, while $X$ ($Z$) syndrome qubits are to extract the syndrome information for quantum error correction. For example, an $X$ ($Z$) syndrome qubit is to measure the stabilizer $\otimes_{j\in\mathcal{N}}X_{j}$ ($\otimes_{j\in\mathcal{N}}Z_{j}$), where $\mathcal{N}$ is the set of neighbor qubits. The syndrome extract circuits for X and Z stabilizers are shown in Fig. 3(b) and 3(c), respectively. In a QEC cycle, the X and Z stabilizers of the code patch are extracted, which will be stored and sent to decoders for error correction (Dennis et al., 2002; Fowler et al., 2012a, d, c; Higgott and Gidney, 2025).
2.7. Pauli-Based Computing Transpilation
In the recent development of QEC, especially in superconducting-circuit-based surface code constructions, the limited connectivity between logical patches has led to the proposal of implementing CX gates via lattice surgery. This approach enables straightforward multi-qubit Pauli measurements at the logical level and supports T-gate implementation through multi-patch Pauli measurements combined with prepared magic states.
Pauli-based computing (PBC) is a quantum computational scheme where qubits undergo a sequence of nondestructive eigenvalue measurements on Hermitian Pauli operators—which may be chosen adaptively based on prior measurement outcomes—and the final result is a classical bit derived from polynomial-time processing of the recorded eigenvalues (Bravyi et al., 2016). PBC can readily describe the lattice-surgery-based surface code computation model. Specifically, in (Litinski, 2019), a method is introduced to remove all Clifford operations in a quantum circuit by (1) commuting the Clifford gates through the T gates and expanding the T gates into multi-qubit Pauli rotations and (2) absorbing the Clifford operations into measurements to form multi-qubit Pauli measurements. The transformation rule can be derived as
(6) $\left\{\begin{array}[]{l l}P\rightarrow P,&\text{if }[P,P_{c}]=0\\ P\rightarrow iP_{c}P,&\text{if }[P,P_{c}]\neq 0\\ \end{array}\right.,$
where $P$ is the Pauli-$\pi/4$ rotation axis (represented as a Pauli operator) or the measurement basis, and $P_{c}$ is the Pauli operator corresponding to the Clifford rotation angles (Litinski, 2019).
2.8. Pauli Grouping
In quantum algorithms, extracting information from the quantum states prepared by a circuit is crucial, especially in applications such as quantum chemistry, where variational quantum eigensolvers require measuring the expectation value of a Hamiltonian (Peruzzo et al., 2014; Tilly et al., 2022; Cerezo et al., 2021). However, quantum hardware natively supports only Pauli-$Z$ measurements on individual qubits. Consequently, evaluating the expectation value of problem Hamiltonians requires decomposing them into Pauli strings that can be measured with existing hardware.
To minimize the number of circuit executions shots, a common strategy involves grouping commuting Pauli strings together to measure them simultaneously (Yen et al., 2023). Two popular approaches are Qubit-wise commutation (QWC) (Verteletskyi et al., 2020) and Group-wise commutation (GC) (Crawford et al., 2021). Suppose we have two $n$-qubit Pauli strings $P_{1}=\bigotimes_{j=1}^{n}\sigma_{k_{j}}^{(j)}$ and $P_{2}=\bigotimes_{j=1}^{n}\tau_{l_{j}}^{(j)}$, where $j$ is the index for qubit, $\sigma_{k_{j}}^{(j)},\tau_{l_{j}}^{(j)}\in\{I,X,Y,Z\}$. Under QWC, any two Pauli strings in the same set must satisfy
(7) $\forall j,\ [\sigma_{k_{j}}^{(j)},\tau_{l_{j}}^{(j)}]=0,$
whereas GC only requires $[P_{1},P_{2}]=0$. Notably, Pauli strings grouped under QWC can be diagonalized using a single layer of single-qubit gates, while GC-based groupings typically require solving for a more complex Clifford circuit to diagonalize the entire set at once.
3. STABSim Design
We provide two different implementations of the stabilizer simulator, a lightweight CPU version and a parallelized CUDA version. The CPU version implements the tableau as a 2-D grid of independent boolean values for simplicity and quickness. Single boolean values in a 2-D array are intuitive to access and manipulate when extending the simulator for tasks beyond direct simulation. CHP operations are performed sequentially, bit-by-bit on the tableau using loops. While straightforward and fast for small simulations, this approach grows significantly at larger qubit sizes. Every Clifford gate update requires a $\theta(n)$ sweep over the tableau to keep the stabilizers up to date, and measurements can be up to $O(n^{2})$ in the worst case. For that reason, we focus on STABSim-GPU for our results.
3.1. GPU implementation
For GPU implementation of STABSim, the stabilizer tableau built in STAB-Sim is stored in global memory and accessed by indexing flattened $x$ and $z$ arrays. Shared memory is avoided for this task since multi-qubit gates and measurements would require reloading and syncing shared memory across blocks. Despite faster memory access in theory, handling memory between shared blocks slowed down circuits with meaningful multi-qubit and measurement gates. Instead, each thread is responsible for its own stabilizer row of the tableau for non-measurement operations. Row-level parallelism avoids race conditions for non-measurement gates and allows the tableau to scale quickly without more operations per thread.
In contrast to a boolean or packed-bit CPU tableau, which must iterate over rows for each gate, STAB-Sim’s GPU design parallelizes row updates, significantly reducing latency for wide or deep circuits. Because a stabilizer tableau contains 2n rows for an n-qubit system, the time to simulate a single gate increases linearly with qubit count in a sequential simulation. However, since stabilizers evolve independently under single-qubit Clifford operations, GPU threads can apply gate transformations to each row concurrently as long as multi-qubit gate dependencies are accounted for.
STAB-GPU also supports $sim2D$, a simulation method which computes all user-defined independent gates at once. This is useful to the extent that the target circuit is limited by the depth of Hadamard, phase, and CNOT gates. However, performance gains depend heavily on circuit structure. For example, applying $sim2D$ to the ancilla initialization and entanglement in a surface code cycle 7 resulted in negligible speedup, likely due to the high frequency of intermediate measurements limiting gate parallelism.
On the other hand, $sim$ should be used for circuits with dynamic measurement operations. Measurement operations particularly comprise the bulk of the simulation time, especially in any circuit where measurements are more frequent than a single end layer. In STAB-GPU, measurement operations are scheduled with block and grid-wide synchronization steps and occasionally utilize atomic operations to collaborate. To start any measurement, the simulator first searches for stabilizer in the tableau that does not commute with a Z-measurement at the target qubit. In practice, this is a search for the first stabilizer row where $x=1$ at the target qubit column. Every thread calculates its own row index, and the first row where $x=1$ is determined by a global $atomicMin$ of all threads after an initial inner-block reduction. If an anti-commuting stabilizer is found, the measurement will be deterministic, and random if not.
In both cases, the outcome and tableau update are computed using the $rowsum$ operation introduced by Aaronson and Gottesman (Aaronson and Gottesman, 2004) for CHP, which is parallelized in STAB-GPU. When the measurement is deterministic, $rowsum$ aggregates all relevant stabilizers into a single scratch row to calculate the measurement result. The number of row operations is worst case $O(n)$ if every stabilizer anti-commutes. Deterministic measurement outcomes rely on stabilizer rows to be aggregated in order, and the inner sum is calculated column by column in $rowsum$. This portion is $\Theta$(n) since every column must participate in the check, so parallelization inside $rowsum$ has high utilization. Each thread calculates its contribution to $rowsum$ based on its unique column bit in the source and target rows. Once that is done, every thread contributes to an atomic $rowsum$, where the final $r$ bit is determined by all of the contributions. For random outcomes, the direction of $rowsum$ is reversed: the selected stabilizer row propagates its update across anti-commuting rows. $rowsum$ is still the dimension being parallelized since its always linear time, but some of the synchronization requirement is removed since the updates to being broadcast to independent stabilizers are independent of each other.
3.2. Scheduling Experiment and Gate-Wise Parallelization
Parallelizing column wise, i.e. assigning other threads to other gates and qubits, effectively shrinks the simulation depth of the circuit, providing another dimension of performance improvement. When parallelizing gate operations, gates should be independent of each other to avoid overwriting tableau data while being used by another thread. Since individual gate operations are so cheap, usually a handful of $xor$ operations, it is easy for threads to jump ahead without synchronization, especially after thousands or millions of operations. This idea lends itself to a scheduler, where multiple independent gates are performed simultaneously. Scheduling at run time seems useful, but did not benefit in our experiments. Even the simplest comparisons used to ensure the next gate could safely proceed take longer than the time to directly compute the current gate in the case of single-pass simulation. By defining flags in the circuit at circuit generation time, it can be preprocessed into chunks safely independent chunks. With $StabGPU$, the user can call $sim(circuit)$ to perform a gate-wise simulation, or $sim2d(cirucit,chunk\_size)$ to parallelize gates within portions of the circuit where operations are independent from one another. In testing QEC-like circuits, little to no advantage was found using gate-parallelization. QEC runtime is dominated by the time to collect mid-circuit measurements, and so improving the small chunks of non-measurement gates in between does not do much for the run time. However, it is included in STAB-Sim because there is no reason why a better suited workload could not take advantage of $sim2d$ to speed up simulation time if the user is interested in very many Clifford gates.
3.3. Pauli Grouping Using STAB-Sim
Pauli grouping is a well used and understood method in quantum chemistry and optimization. Because of this usefulness, and applications to algorithms 3 and 4 in our T-Gate transpiler, we quickly discuss functions for grouping in the context of binarized stabilizers.
Using $x$ and $z$ as the bits of the $X$ and $Z$ components of the stabilizers (Pauli strings), $(x^{i}_{1}\land z^{i}_{2})\oplus(x^{i}_{2}\land z^{i}_{1})=0$ if they commute at each qubit $i$. Running an xor product can extend the commutation check from qubit-wise to group-wise over the full Pauli string:
$\bigoplus_{i=1}^{n}((x^{i}_{1}\land z^{i}_{2})\oplus(x^{i}_{2}\land z^{i}_{1}))=0$
if $P_{1}$ and $P_{2}$ commute. If the binary vectors can be packed into larger data types (e.g. 32, 64-bit words), they can be quickly processed with vectorized bitwise instructions. For group-wise commutation, the xor product reduction can be replaced in favor of a population count on the vectorized results if its an available intrinsic. For example, if
$\mathrm{popcount}\left(\left\{i\in[n]\;\middle|\;[P_{1}^{(i)},P_{2}^{(i)}]\right\}\right)\bmod 2=0,$
then $P_{1}$ and $P_{2}$ group-commute.
To evaluate the commutation of a Pauli string $P$ against an entire stabilizer tableau using matrix operations, the following matrix expression for all stabilizers $S$ in the tableau is used:
$(x_{P}\cdot Z_{S})+(z_{P}\cdot X_{S})\mod 2=0$
if $S$ commutes with the full tableau. Here, $x_{P}$ and $z_{P}$ are the binary row vectors representing the $X$ and $Z$ components of the Pauli string $P$. $X_{S}$ and $Z_{S}$ are the binary matrices corresponding to the $X$ and $Z$ components of all stabilizers $S$ in the tableau. The final sum matrix is an $n\times n$ matrix of the qubit-wise commutation results from $[P,S_{i}]$ where $i$ is each row.
3.4. Clifford Push Through Transpiler
Transpiling a Clifford+T universal quantum circuit into Pauli measurements with a separate T-layers has great potential to make efficient use of quantum machinery by abstracting away Clifford operations, and minimizing expensive magic state operations. Commuting T-layers can be optimized after separation by removing redundant eighth rotations and pushing them to the Clifford layer as quarter rotation gates.
3.4.1. Measurement Tableau and T-Tableau Construction
Algorithm 2 Tableau Construction
1:Initialize an empty tableau $T_{tab}$ to store $pi/4$ rotation gates ($T$ and $T^{\dagger}$), and an $n\times n$ identity tableau $M_{tab}$.
2:for all gates $G$ at qubit $q$ starting from the end do
3:     if $G=T_{q}$ then
4:         Append an $n$-qubit $Z_{q}$-stabilizer to $T_{tab}$ and set $r\leftarrow 0$
5:     else if $G=T^{\dagger}_{q}$ then
6:         Append an $n$-qubit $Z_{q}$-stabilizer to $T_{tab}$ and set $r\leftarrow 1$
7:     else if $G$ is Clifford then
8:         Apply $G$ to $M_{tab}$ and $T_{tab}$ using CHP rules
9:     end if
10:end for
With the target Clifford+T circuit arranged in time order, gates are extracted from the end and added to one of two sets of tableaus. The first tableau is an n-qubit identity tableau of only stabilizers, used for tracking Clifford operations absorbed into the measurement. The second is an empty n-qubit tableau used for tracking the commutation rules of T gates as they are pushed through the circuit. If the gate extracted is a T gate, it is appended to the end of the T tableau as an n-qubit Pauli string with a Z stabilizer at the qubit index of the T gate. If the gate extracted is a Clifford operation, it is applied to the measurement (M) tableau by CHP gate operations. It is also applied to the T tableau by the same rules, but only acts on the T gates which have been appended up to that point. For example, if from the end of the original circuit, no T gates have been extracted yet, then the Clifford gate will have no affect on the T tableau as there are no interactions with the T gates that come before it in time order.
Figure 4. Gate-by-gate extraction from the reversed circuit to construct the Measurement and T Tableaus. The Measurement tableau is a standard CHP tableau. The T-tableau treats T-gates as Z-stabilizers at the T-gate target qubit. Clifford gates are applied to any stabilizer that has been appended at the point where the gate occurs in reverse-time order.
3.4.2. T-Layer Separation
Algorithm 3 T-Separation
1:Initialize tableau list $\mathcal{P}\leftarrow[\ ]$
2:Create tableau $P_{0}$ and append the last stabilizer in $T_{\text{tab}}$
3:Append $P_{0}$ to $\mathcal{P}$
4:for all stabilizers $S$ from second-to-last to first in $T_{\text{tab}}$ do
5:     $appended\leftarrow\textbf{false}$
6:     for all $P\in\mathcal{P}$ starting from $P_{0}$ do
7:         if $[S,P]=0$ then
8:              Append $S$ to $P$
9:              $appended\leftarrow\textbf{true}$
10:              break
11:         end if
12:     end for
13:     if not $appended$ then
14:         Create new tableau $P_{\lvert\mathcal{P}\rvert}$ and append $S$
15:         Append $P_{\lvert\mathcal{P}\rvert}$ to $\mathcal{P}$
16:     end if
17:end for
Once the full circuit has been parsed and M and T tableaus have been created, the T tableau is split into commuting groups as follows. Starting from the end of the T tableau (the last appended stabilizer, i.e. the first T gate in time order) a new empty tableau that we will call $P_{0}$ is made and the stabilizer is appended. Each subsequent stabilizer from the end of the T tableau is checked against the new P tableau using the commutation method discussed in  3.3. If the subsequent stabilizer commutes with each Pauli string in the first P tableau, it is appended to that tableau. If there is an anti-commutation between the stabilizer being checked and a stabilizer within the tableau, a new tableau ($P_{1}$) is created. For the next stabilizer from the end of T, it is first checked against $P_{0}$. If it commutes with $P_{0}$ it is appended to $P_{0}$. If not, the stabilizer is checked against $P_{1}$ and appended to $P_{1}$ if it commutes. If it anti-commutes, the stabilizer starts a new tableau $P_{2}$. This process continues from the end of T for all stabilizers in T until T is empty. After this process is complete, there will be some number of self-commuting P tableaus (at least one, at most the number of T gates in the original circuit).
Figure 5. T-Separation. The T-Gate tableau is separated into commuting groups by placing the first T-stabilizer at the end of the original circuit order. If the stabilizer does not commute, it cannot be grouped into the same T-layer, and must be appended to an earlier T-Layer or create its own.
At this point, the basic measurement circuit and T layers have been created. The total number of T-stabilizers is counted to later determine when optimization is done. Clifford rotations are now extracted from the P tableaus with the goal of reducing the depth of the T layers. To do this, repeated stabilizers within each tableau are counted, starting from the last P tableau. The P tableaus are in reverse time order, meaning the first entry of $P_{0}$ contains the very last T gate in the original circuit. We want to push Clifford gates to the end, towards the measurement circuit, to absorb any leftover Clifford rotation. For every two occurrences of a stabilizer, a quarter rotation gate can be extracted. Likewise, for every four occurrences of a quarter rotation gate, a full rotation is made, and all eight T stabilizers can be removed without pushing any operations through to the measurement tableau.
3.4.3. T-Layer Optimization
Algorithm 4 T-Optimization
Count initial total number of rows in all $P_{i}\in\mathcal{P}$, store as $N_{\text{prev}}$
Initialize convergence flag $converged\leftarrow\textbf{false}$
while not $converged$ do
     for $i=\lvert\mathcal{P}\rvert-1$ down to 00 do
         while there exist two identical stabilizers $S$ in $P_{i}$ do
              Remove both copies of $S$ from $P_{i}$
              Append one copy of $S$ to $P_{i+1}$
              for all stabilizers $S_{T}$ in $P_{i+1}$ do
                  if $[S_{T},S]\neq 0$ then
                       Apply $\text{rowsum}(T,S)$
                  end if
              end for
              Remove $S$ from $P_{i+1}$ and append to $P_{i+2}$
         end while
     end for
     Count the total number of rows in all $P_{i}\in\mathcal{P}$, store as $N_{\text{new}}$
     if $N_{\text{new}}=N_{\text{prev}}$ then
         $converged\leftarrow\textbf{true}$
     else
         $N_{\text{prev}}\leftarrow N_{\text{new}}$
     end if
end while
$\lvert\mathcal{P}\rvert$ is the number of $\pi/4$ rotation gate layers after minimization. The sum of all non-Identity Pauli operators in all $P_{i}$ stabilizers is the number of $\pi/4$ rotation gates.
We start by rightward gate pushing from the last P tableau, i.e., the first T gates in original circuit order. Repeat stabilizers in the tableau are counted, removed from the tableau, and the corresponding quarter rotations are then checked For the second to last P tableau, repeat stabilizers are once again counted and removed, but now need to be commuted through the P tableau to the right before being added to the beginning of the measurement circuit. To do this, one copy of the repeated stabilizer is appended to the next P tableau and checked to see if it commutes with every stabilizer in the tableau. If it does, it is passed through to the next tableau. If not, a $rowsum$ 1 is called on the appended stabilizer with each row of the tableau that does not commute. $rowsum$ accumulates commutation the effect of an anti-commuting Pauli string throughout the stabilizer tableau by summing the contributions of each stabilizer Pauli and corresponding phase bit $r$ (Aaronson and Gottesman, 2004). The sum modulo 4 of the appended Pauli string respective non-commuting stabilizer becomes the new phase bit $r$ of that stabilizer. $rowsum$ is originally used in the CHP formalism to propagate the effect of measurement on the tableau. To satisfy the commutation rules for pushing a Clifford rotation gate through another arbitrary rotation (Litinski, 2019), a phase of $i$ is added to the $rowsum$ product.
Figure 6. T Optimization. Every two repeating stabilizers are removed from the tableau and pushed left, following the $rowsum+i$ routine for each unique Pauli string.
Once a pass of Algorithm 4 is complete, the total T-stabilizers remaining are counted. If no repeat stabilizers were extracted during optimization, the process ends. Otherwise, optimization is repeated to extract new repeat stabilizers created by $rowsum+i$, until the total number of T-stabilizers no longer changes. After optimization, the $P$ tableaus can be recombined to a larger $T$ tableau. The stabilizers of the $T$ tableau are the self-entangled layers of the $\pi/4$ rotation gates that remain. In Figure 2, every non-Identity Pauli operator in all the layers is counted as a T-rotation, for comparison with the number states needed for T-gates in the original circuit.
4. Evaluation
CPU simulator evaluations were all performed on an AMD EPYC 7502 compute node with a 3.35 GHz boost clock and 256 GB of RAM. GPU evaluations were performed on an NVIDIA H100 SXM node. These platforms were chosen for a stable and consistent test environment versus a laptop or desktop that may have power saving modes and background processes. Tests for STABSim-CPU and STABSim-GPU were compiled with GCC 10.3.0, and STABSim-GPU used CUDA 12.1.
4.1. QEC Simulation
Stabilizer simulation is made feasible in time and size by the relatively slow growth of the stabilizer tableau as qubit count increases compared to other simulation methods. However, measurement operations are $O(n^{2})$ worst-case in the original CHP formalism, one dimension due to every stabilizer potentially being checked, and another from every column participating in $rowsum$ when the row check conditions are met. Measurements greatly bottleneck the time for dynamic circuits, especially QEC codes where many repeated stabilizer measurements are the primary interest for evaluation. Stim improves on the deterministic measurement worst case, and, combined with 256-bit SIMD instructions, we find that Stim outperforms STAB-GPU in surface code up to distance 33 with 2179 qubits/tableau columns. Beyond distance 33, STAB-GPU scales better than existing stabilizer simulators due to the lower extra cost associated with adding a qubit to the simulation 7. The main constraint in STAB-GPU is the synchronization necessary for correctness, especially in measurement, which leads to a closer-to-linear increase in simulation time as a function of circuit size. STAB-CPU is implemented as a boolean tableau without bit-packing or other methods to speed up wide circuits. This allows STAB-CPU to be very lightweight and extensible. Startup time for STAB-CPU is less noticeable than STAB-GPU, and it has the edge in small simulations where the overhead for launching a CUDA kernel has a meaningful impact on simulation time.
Figure 7. Full single-shot tableau simulation
of planar surface codes as code distance increases. The structure follows the example given in 2.4.
Gate-wise parallelization was also tested for a planar surface code. In that case, we found that even in very large tableaus, simulation time remained almost the exact same due to the time-dependence of measurement operations. Time improvement gained by parallelized qubit initialization and entanglement between many data and ancilla qubits was minimal compared to the impact of measurement. Figure 7 shows how different simulators scale in a less structured workload that still mimics QEC. In this scenario, half of the qubits have a Hadamard or phase gate applied at random. The initialized qubits in the first half are used as the control for a CX gate to each corresponding qubit in the second half. 20% of the second half qubits are then measured, chosen at random. This cycle is layered $log_{2}(n)$ times for the sake of having a non-trivially deep circuit while keeping reasonable testing times, with new randomly chosen values for each layer.
Figures 7 and  8 display similar scaling trends, with a tradeoff from Stim to STAB-GPU at $\approx 2000$ qubits and around the 10 second mark for simulation time. Qiskit-Aer method=’stabilizer’ stays closer to Stim in the random test, in part because Stim’s time complexity improvement in deterministic measurements is utilized for only about half of the measurements, versus nearly all measurements in surface code. STABSim-CPU has similar scaling concerns as Stim, but performs well on the compute node to a surprising extent. Stim should perform far fewer row operations per Clifford gate than STABSim-CPU due to the packed 256-bit AVX instructions it uses. One possibility is that frequent measurement operations over thousands of qubits require enough bit masking operations to outweigh the time saved on simple Clifford gates. One benefit the testing platform has over a standard laptop CPU is a large 128 MB L3 cache. Fewer cache misses could potentially lead to less gain for Stim from an efficient data layout, even in very large tableaus.
Figure 8. Full single-shot tableau simulation
of partially random circuit. Half of the qubits are initialized with Hadamard or phase gates at random, and control a corresponding qubit in the second half. $20\%$ of the second half qubits are measured. The process is repeated $log_{2}(n)$ times with new random gate applications.
4.2. Pauli Grouping
To demonstrate the utility of STABSim, we apply STABSim to Pauli grouping problems. We focus on grouping the Pauli strings from molecular Hamiltonians, which are widely used in quantum chemistry applications. The Hamiltonians were chosen for varying degrees of freedom and measurement gate sizes, while remaining practical to simulate with our state vector simulator. The detailed information about the molecular configuration settings and the Pauli group terms are shown in Table 2.
To group the Pauli strings, the following steps were employed. First, Pauli terms are sorted up front by the weight of the Pauli term coefficients. Groups of commuting Pauli terms are then greedily constructed starting with the highest weight terms by either QWC or GC. This results in higher weight terms being grouped earlier so sampling resources can be prioritized to get a higher fidelity from more shots on larger terms. Table 2 counts the total number of grouped gates. To verify the correctness of the ground state energy result after grouping, all molecular Hamiltonians were evaluated with a statevector simulator we have developed.
Our tableau-based simulator enables fast and scalable implementation of both strategies with xor product commutation checks. The time spent performing grouping and generating measurement settings was negligible compared to the overall simulation time, especially when considering the time for statevector-based energy verification. VQE in particular still faces challenges in exponential post-processing complexity (Scriva et al., 2024), but Pauli grouping is a useful and scalable pre-processing step that can be used with any Pauli-term based quantum algorithm to reduce resource overhead.
Table 2. Gates required for evaluation after grouping small molecular Hamiltonians. Initial basis gates, Qubit-Wise, and Group-Wise gates are shown, along with the % improvement from QWC to GWC in evaluation gate count.
Molecule Orbitals Initial QWC GWC % Change
H4 4 3428 717 275 61.65%
LiH 3 756 449 316 29.62%
LiH 4 3428 878 427 51.37%
H6 18 16196 3844 1858 51.66%
H6 22 16196 3850 1946 49.45%
BeH2 6 16196 1681 749 55.44%
BeH2 7 27124 3627 1580 56.44%
4.3. PBC Transpilation
Like Pauli grouping, the T-Transpiler application relies on fast commutation checks to be efficient. Figure 9 shows the time to perform Algorithm 3 to unpack an example circuit, and the time to perform optimization Algorithm 4, including the greedy commutation grouping described in Section 3.4.3. Similar to the Pauli grouping use-case, the purpose of the T-transpiler is to reduce quantum resources. Instead of reducing rotations for measurement settings, the quantity being optimized is the number of T-states in a static circuit along with trivializing the Clifford gates as rotations of the measurement basis.
Figure 9. Time to process and optimize the circuit for select circuits using Python to track commutations as boolean values, versus STABSim.
Circuit Python vs Clifford+T T- Optimized T-Ratio
STAB Time Gates Gates Gates
Adder N10 14.26x 189 88 16 2.0x
Adder N4 34.61x 25 4 4 1.0x
Adder N433 2.74x 6196 1536 768 2.0x
Basis Test N4 51.88x 577 156 132 1.18x
Basis Trotter N4 16.64x 9819 1860 1578 1.18x
Bell N4 17.19x 386 70 68 1.03x
Dnn N16 8.17x 30016 6656 5920 1.12x
Dnn N2 46.46x 3414 760 672 1.13x
Dnn N8 16.56x 15008 3328 2960 1.12x
Fredkin N3 19.28x 22 4 4 1.0x
Gcm H6 7.97x 17719 4852 4486 1.08x
Inverseqft N4 13.29x 99 28 28 1.0x
Ising N10 30.14x 3152 974 928 1.05x
Ising N26 43.93x 1116 332 324 1.02x
Knn N25 11.79x 816 204 192 1.06x
Multiplier N15 3.32x 454 144 42 3.43x
Multiplier N45 6.77x 4574 1512 396 3.82x
Multiply N13 11.67x 92 24 14 1.71x
Pea N5 13.69x 523 166 166 1.0x
Qaoa N3 28.38x 53 8 8 1.0x
Qaoa N6 28.65x 2928 672 544 1.24x
Qec En N5 27.47x 25 1 1 1.0x
Qf21 N15 18.42x 646 170 162 1.05x
Qft N18 19.05x 2150 596 596 1.0x
Qft N18 Iter1 29.37x 3290 1677 1333 1.26x
Qft N18 Iter2 20.16x 27292 14383 11463 1.25x
Qft N4 38.83x 310 92 92 1.0x
Qpe N9 53.14x 334 90 86 1.05x
Qram N20 32.09x 276 80 48 1.67x
Sat N7 32.62x 213 40 24 1.67x
Seca N11 29.27x 200 32 16 2.0x
Simon N6 21.12x 56 8 8 1.0x
Swap Test N115 10.99x 2866 736 680 1.08x
Swap Test N25 13.76x 592 152 140 1.09x
Teleportation N3 56.61x 8 1 1 1.0x
Toffoli N3 22.36x 20 3 3 1.0x
Variational N4 51.23x 363 100 92 1.09x
Vqe N4 28.66x 639 188 186 1.01x
Vqe Uccsd N4 15.41x 952 160 160 1.0x
Vqe Uccsd N8 4.48x 33969 3816 3742 1.02x
Wstate N27 26.17x 1704 416 416 1.0x
Wstate N3 71.54x 67 14 14 1.0x
Wstate N76 3.42x 2488 416 416 1.0x
Table 3. Clifford+T transpilation speedup using STABSim. The total Clifford+T gates are counted for comparison after optimization. T-ratio shows the reduction of one-eighth rotations from the initial number of T-type gates.
Most of the substantial time gain found with STAB-Sim is an improvement in the optimization step. Processing is primarily an issue of circuit size, and does not require commutation checks or other applications that can be sped up by bitwise methods. Table 3 shows how static Clifford+T circuits change under Clifford absorption into measurement and T-count optimization. The T-stabilizers remaining after optimization comprise the entire remaining circuit, apart from the now-merged and rotated measurement operations at the end. Each stabilizer in the T-tableau after the full processing is an entangled layer of $\pi/4$ rotational gates in the X, Y, or Z basis, given by the Pauli operator on each qubit. A phase bit of 1 in the stabilizer means the basis rotation in that layer is negative. Interestingly, circuits that showed the least potential for T-State reduction in Table 3 had the best improvement in checking the reduction from STAB-Sim versus our Python reference, as seen in Figure 9. The circuits with more trivial T-States tended to converge quickly and eliminate many T-layers, which reduced the size and frequency of loops in Algorithm 4, where commutation checks become a bottleneck. The difference in small circuits can also likely be attributed to the compiled nature of C++, especially the four qubit adder, where the only work is to add up the Clifford rotations for the end of circuit measurement settings, and check if there is any entanglement between the minimal four T-gates.
In our testing, the QFT circuits compiled into Clifford+T circuits by applying the Solovay-Kitaev algorithm (Dawson and Nielsen, 2006) had the fewest trivial T-gates, meaning they were often barred from combining with other T-rotations. The compiled multiplier circuits tended to eliminate T-roations quickly, and mostly converge after the first round of optimization.
Similar to the optimization performed in Table 2, the T-transpiler has application where simulation resources are limited or fewer magic states speed up circuit run times on a quantum device. If T-gates are sufficiently eliminated from the circuit, there is also a lower barrier towards Clifford-dominated circuits with few T-states being simulated by stabilizer simulation, as described by Bravyi and Gosset (Bravyi and Gosset, 2016).
5. Related Work
Many projects have been developed on the concept of simulating a quantum system from its stabilizers since CHP was proposed by Aaronson and Gottesman (Aaronson and Gottesman, 2004). Cirq, Qiskit, and other projects all implement stabilizer simulators as functions in their libraries (Developers, 2021; Qiskit Development Team, 2021). More recently, the Stim project made large gains in QEC sampling via Pauli frame simulation after the initial stabilizer simulation is complete (Gidney, 2021).
Cuda-QEC leverages GPU for accelerated decoding of error syndromes (NVIDIA Corporation & Affiliates, [n. d.]). However, it does not provide a simulation method for evaluating error correcting codes. Qua-SARQ performs GPU accelerated simulation of Clifford gates on a stabilizer tableau, to replace the work of Stim in the equivalence checking applications (Osama et al., 2025). Crucially, Qua-SARQ only simulates Clifford gates, which, given the similarity to large binary matrix operations, are a portion of stabilizer simulation that end up being straightforward to parallelize. Qua-SARQ does not address measurement gates, which turn out to be the most costly portion of simulation and the crux of syndrome extraction in QEC.
Pauli measurements and Pauli-Based Computation (PBC) are discussed significantly by Litinski in the context of surface code operations for fault-tolerance (Litinski, 2019). Peres and Galvão recently expanded on PBC compilation with a standalone Python compiler, and space/time tradeoff using auxiliary qubits (Peres and Galvão, 2023). To our knowledge, our work is the first to formulate PBC in the context of a stabilizer tableau, where highly efficient bitwise stabilizer operations can be exploited for compilation. TQEC
The stabilizer tableau in this work also provides a natural platform to simulate the rotated Pauli measurements that are pushed through the circuit. Combined with our statevector simulator used in Section 4.2, or using a few T-gate method similar to what was proposed by Bravyi and Gosset (Bravyi and Gosset, 2016), the full compiled Clifford+T circuit can be evaluated by simulation.
6. Conclusion
Large stabilizer simulations become more and more of interest as quantum computing moves towards intermediate scale and eventual fault tolerance. Applications benefiting from large scale stabilizer-based computing can be found in higher-distance error correcting codes, quantum verification, limited Clifford+T circuits, heterogeneous codes, and more (Bravyi et al., 2024b; Fowler et al., 2012b; Bravyi and Gosset, 2016; Stein et al., 2024). STAB-Sim shows promise in evaluating growing QEC codes or other large circuits in reasonable amounts of time by scaling better than any existing stabilizer simulator. We furthermore show how the extended capabilities of STAB-Sim directly benefit other domains beyond direct simulation.
Work in these areas has also inspired ideas for future directions. For example, while tracking Clifford and T push-through is relatively straightforward for static circuits using a limited amount of tableaus, it would be interesting to investigate how effectively a dynamic circuit with mid-circuit measurements could be transpiled, and if such a transpiler would be beneficial. Application of our fast Clifford simulatior to large Clifford+T circuits using Bravyi and Gosset’s algorithm would also be interesting to explore what universal quantum circuits could potentially be simulated in practical time.
Acknowledgment
This research was supported by the Quantum Algorithms and Architecture for Domain Science Initiative (QuAADS), under the Laboratory Directed Research and Development (LDRD) Program at Pacific Northwest National Laboratory (PNNL). M. Wang and A. Li were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (QSC). This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830.

### v2

@@
 Abstract.
-Efficient stabilizer-based computation is a key task in quantum computing as it provides a classically tractable insight into quantum algorithms, which otherwise quickly become impossible at scale on a classical device.
-Formulations in quantum error correction, hardware verification, hybrid computing, and more, all benefit from the ability of Clifford gates to be accurately evaluated by classical computers.
-However, current stabilizer simulators have significant limitations in scaling and use-case beyond direct simulation.
-To address this, we present a GPU-accelerated tableau stabilizer simulator that scales efficiently in direct circuit simulation, and supports a range of stabilizer-based tasks.
-We show how the simulator is used to quickly calculate Pauli commutation groupings between Pauli strings to reduce the number of circuit runs in quantum chemistry problems.
-We also present a Clifford+T circuit transpiler based on STABSim, which uses the simulator framework to track relations in non-Clifford rotations as Clifford gates are absorbed into the measurement step of the circuit.
-The T rotation layers left behind are then further optimized in the simulator structure to create a reduced circuit containing only the minimum non-Clifford rotations, improving the time to simulate and reducing the cost to run on real quantum devices.
+The quantum stabilizer formalism became foundational for understanding error correction soon after the realization of the first useful quantum error correction codes.
+Stabilizers provide a way to describe sets of quantum states which are valid codewords within a quantum error correction (QEC) scheme.
+Existing stabilizer simulators are single threaded applications used to sample larger codes than is possible with other methods.
+However, there is an outstanding gap in the scaling and accuracy of current simulators for QEC as quantum computing exceeds hundreds of qubits, along with an under-utilization of the capabilities of highly-efficient stabilizer simulation across other quantum domains.
+In this work, we present the first GPU-accelerated tableau stabilizer simulator to scale better than CPU methods in QEC workloads, by trivializing Clifford gates and exploiting the large parallelism of dedicated GPUs with CUDA warp-level primitives to quickly overcome costly measurement gates.
+We then implement a new error model that captures non-unitarity in T1/T2 error channels much faster and with exact accuracy for most physical qubits, demonstrate a chemistry use case, and present a new Clifford+T to Pauli-Based Computing (PBC) transpilation optimization through our simulator.
 copyright: none
 1. Introduction
-Quantum computing holds the promise of dramatically accelerating solutions to classically hard problems in areas such as cryptography (Shor, 1999; Gisin et al., 2002), optimization (Farhi et al., 2014; Han and Kim, 2002), quantum chemistry (Georgescu et al., 2014; Kandala et al., 2017), and finance (Rebentrost et al., 2018; Woerner and Egger, 2019), etc.
+Quantum computing holds the promise of dramatically accelerating solutions to classically hard problems in areas such as cryptography (Shor, 1999; Gisin et al., 2002), optimization (Farhi et al., 2014; Han and Kim, 2002), quantum chemistry (Georgescu et al., 2014; Kandala et al., 2017), etc.
 In recent years, hardware demonstrations have shown an increasing number of qubits while reducing error rates.
 Quantum devices capable of demonstrating of quantum advantage in certain tasks have been reported (Arute et al., 2019; Wu et al., 2021; Gao et al., 2025).
-However, to gain the full potential of quantum computing, the error of quantum systems needs to be carefully managed and corrected (Nielsen and Chuang, 2010) using quantum error correction (QEC) codes.
-QEC encodes dozens of physical qubits into one logical qubit and actively detecting and correcting errors that occur in the quantum systems.
+However, to gain the full potential of quantum computing, the error of quantum systems needs to be carefully managed and corrected (Nielsen and Chuang, 2010) using QEC codes.
+QEC encodes many physical qubits into one logical qubit and actively detects and corrects errors that occur in the quantum systems.
 QEC lays the foundation for fault-tolerant quantum computing.
 Early demonstrations of small-scale QEC codes have shown encouraging results (Ryan-Anderson et al., 2021; Krinner et al., 2022; Acharya et al., 2023; Bluvstein et al., 2024; Acharya et al., 2025), but the path toward mid-term and far-term devices featuring thousands of logical qubits demands significant advances in architecture, control, and simulation (Gidney and Ekerå, 2021).
-In particular, designing and verifying different QEC codes such as surface codes, color codes, and especially understanding the error threshold and the noise performance of logical gate operations, requires efficient large-scale classical simulation tools.
+In particular, designing different QEC schemes while understanding the error threshold and noise performance of logical gate operations, requires very efficient classical simulation tools.
 Figure 1. Simulation time of surface codes with distances ranging from 3 to 71, using Qiskit method=’stabilizer’ and Stim on an AMD EPYC 7763 processor.
-QEC codes primarily consist of Clifford operations, which can be efficiently modeled using the stabilizer formalism (Gottesman, 1998; Nielsen and Chuang, 2010).
+QEC codes primarily consist of Clifford operations and measurements, which can be efficiently modeled using the stabilizer formalism (Gottesman, 1998; Nielsen and Chuang, 2010).
 Unlike general quantum circuit simulators that scale exponentially with qubit count, stabilizer simulators scale polynomially, enabling large-scale QEC design verification.
-Despite this efficiency, existing stabilizer simulators, including Qiskit (Qiskit Development Team, 2021) and Stim (Gidney, 2021).
 Qiskit, in addition to its widely used quantum compilation and emulation features, offers a stabilizer simulation backend that can simulate noisy Clifford circuits if all errors in the noise model are also Clifford errors.
@@
 Both Qiskit and Stim highlight the value of specialized stabilizer simulation in quantum computing research, especially for QEC protocols.
-However, as we move forward from the Early fault-tolerant era to Megaquop machines (Preskill, 2018; Katabarwa et al., 2024; Preskill, 2025), we expect to expand the simulation probabilities of stabilizer simulators, targeting large-scale multi-logical-qubit code block simulations.
-However, Stim’s performance often becomes limited by CPU-centric design constraints such as memory bandwidth and the relatively low degree of parallelism available in general-purpose processors.
-As shown in Figure 1, while Stim performs significantly better than Qiskit, both Qiskit and Stim scale poorly to larger QEC codes.
-Large surface-code circuits with repeated rounds of syndrome extraction, fast feedback, and intricate multi-qubit operations—like those in lattice surgery—push the limits of CPU-only approaches.
-In particular, the sheer volume of stabilizer updates, measurements, and state transformations can quickly overwhelm the available computing resources, resulting in long simulation times and restricting the scale of feasible experiments.
-In addition, as Stim is optimized towards QEC applications, other applications relying on Clifford and Pauli operations can have difficulties being mapped onto the framework to leverage the power of Stim.
+However, as we move forward from the early fault-tolerant era to Megaquop machines (Preskill, 2018; Katabarwa et al., 2024; Preskill, 2025), we expect to expand the simulation capabilities of stabilizer simulators, targeting large-scale multi-logical-qubit code block simulations, with individual patch and multipatch code distances from tens to hundreds.
+In this regime, the polynomial overhead becomes a challenge for state-of-the-art simulators.
+Large QEC circuits with repeated rounds of syndrome extraction, fast feedback, and many qubit operations push the limits of CPU-only approaches.
+Operations like lattice surgery are key for fault tolerance, and have modern tools such as TQEC to enable the construction of enormous fault-tolerant pipelines (noa, 2025).
+However, CPU-centric design constraints such as memory bandwidth and low parallelism, place a practical time constraint on actually characterizing errors in these huge architectural constructions.
+As shown in Figure 1, while Stim is significantly better than Qiskit, both Qiskit and Stim scale poorly to larger QEC codes.
+Furthermore, to take advantage of Stim’s efficiency via Pauli frames, physical qubit errors must be approximated using Pauli operations, which is an assumption that quickly diverges from reality when considering common non-unitary errors such as qubit relaxation.
+Particularly, the volume of stabilizer updates across the stabilizer tableau and scaling faster per-qubit measurement costs quickly overwhelm sequential CPU pipelines, resulting in long simulation times and restricting the scale of practical investigation.
+In addition, as Stim is highly optimized for shot-based QEC evaluation, other applications relying on Clifford and Pauli operations are difficult to integrate.
 To tackle these challenges, we propose a GPU-accelerated stabilizer simulator that exploits the massive parallelism and high-bandwidth memory of modern Graphics Processing Units (GPUs).
-Originally designed for graphics workloads, GPUs today provide thousands of concurrent threads and efficient memory subsystems, making them interesting for the large matrix-like operations at the core of QEC simulation.
-Previous efforts on GPU implementation did not yield performance improvement (Gidney, 2021), citing the low arithmetic intensity of XOR operations at the core of stabilizer simulation.
-CPUs certainty outperform GPUs for small and sequential arithmetic by design.
-However, dedicated GPUs have their own bitwise and synchronization capabilities that can safely enable many operations at once, even in the time-order sensitive context of quantum circuit simulation.
-Furthermore, comparisons and longer arithmetic operations appear frequently during measurement operations, which are the bottleneck of QEC simulation times.
-Our approach encodes stabilizer states directly on the GPU, orchestrating operations so that they can be applied in parallel, thereby significantly reducing simulation time and time scaling for larger circuits.
-This design enables studies of mid-term and far-term QEC protocols, including high-distance surface codes and complex lattice surgery which scales impractically on lightly-threaded CPUs.
-We still provide lightweight CPU version which performs within the realm of other modern stabilizer simulators.
-Both versions are flexible with fast stabilizer and Pauli-based computing features beyond direct simulation.
-As we will show, efficient stabilizer simulation has useful applications for fields beyond QEC, including chemistry simulation and circuit optimization.
-This work will detail the following contributions:
+Originally designed for graphics workloads, GPUs today provide thousands of concurrent threads and efficient memory subsystems, making them an interesting platform for the large matrix-like operations at the core of QEC simulation.
+Previous efforts on GPU implementation either did not yield performance improvement (Gidney, 2021), or could not perform measurement, which is the key operation to QEC (Osama et al., 2025).
+Stim cited the low arithmetic intensity of bit operations at the core of stabilizer simulation as the reason for GPU under-performance.
+For measurement, the strict time-order sensitivity and cost of branching posed a barrier that has not been addressed until this work.
+Our approach encodes stabilizer states directly on the GPU, orchestrating operations to be applied in parallel with minimal synchronization.
+Where time-order synchronization is absolutely necessary, extremely efficient CUDA warp-level primitives and optimized block-wide cooperations are employed, allowing us to rapidly clear synchronization and branching bottlenecks.
+These techniques greatly reduce simulation time as the qubit number and circuit depth increase.
+This design enables studies of large-scale QEC protocols, including high-distance QEC codes and complex lattice surgery operations which scale impractically on lightly-threaded CPUs.
+We still provide a CPU version comparable with existing stabilizer simulators as a flexible platform for validating new stabilizer-based algorithms.
+Both versions support and have been used to develop Pauli-based computing features well beyond direct simulation.
+As we will show, efficient stabilizer simulation is necessary for large sale QEC, but also has surprising applications for better qubit noise modeling, chemistry simulation, and quantum circuit optimization.
+This work thus detail four core contributions:
   1. (1)
-    The first quantum stabilizer simulator to show performance gains via GPU in Clifford+Measurement circuits.
+    The first quantum stabilizer simulator to show a performance gain via GPU in Clifford+Measurement simulation.
+Within a hundreds of qubits, STABSim exceeds state-of-the-art simulators in QEC experiments thanks to highly-efficient CUDA warp primitives combined with careful synchronization for accuracy.
   2. (2)
-    Improved scaling among stabilizer simulators as qubit size increases, lowering the time to evaluate a full surface code simulation and measurement after $\sim 2000$ qubits.
+    Implementation of a new qubit noise model that can be sampled with much more efficiency and accuracy than existing simulators.
+When $T1$ coherence equals or exceeds $T2$, where most physical qubits reside, our implementation only requires tableau sampling with constant overhead, but has exact accuracy to the physical noise channel.
+Even when $T2$ is greater, our implementation still requires much fewer tableau samples than other exact methods.
   3. (3)
-    Integrated Pauli-Based Computing (PBC) capabilities and a direct quantum chemistry use-case shown for VQE.
+    A quantum chemistry use-case for VQE using fast bitwise Pauli string methods available on STABSim.
   4. (4)
-    A novel Clifford+T to PBC transpiler using the stabilizer simulator, showing how stabilizer simulation can be leveraged to improve non-Clifford circuits by reducing T-gate state count by 35-70%.
-This transpiler is the first to our knowledge to automatically convert any static Clifford+T circuit to a PBC circuit using bitwise methods, and the first to perform T-state optimization.
+    A novel Clifford+T to PBC transpiler using the stabilizer framework.
+We detail how stabilizer simulation can improve universal circuit architectures by optimizing T-gate count, and show how our method greatly improves the speed of transpilation versus other frameworks.
 2. Background
@@
 At the core of most quantum computing circuits are Pauli operators, which consist of tensor products of Pauli operators $X$, $Y$, $Z$, and identity operator $I$.
-With the combination of phase factors {±$\{\pm 1,\pm i\}$, Pauli operators form Pauli groups across $n$ qubits.
+Pauli operators, combined with phase factors $\{\pm 1,\pm i\}$, form Pauli groups across $n$ qubits.
 The Clifford group comprises all unitary operators that map Pauli operators to other Pauli operators under conjugation.
-In practice, this group is generated by a small set of gates, e.g., single-qubit Hadamard gates $H$, phase gate $S$, and $CX$ gates, represented in computational basis as
-(1) $H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix},\,S=\begin{pmatrix}1&0\\ 0&i\end{pmatrix},\,\mathrm{CX}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$
 Proved by the Gottesman-Knill theorem, quantum circuits restricted to Clifford gates and Pauli measurements can be simulated on a classical computer in polynomial time with respect to the circuit size and number of qubits (Aaronson and Gottesman, 2004; Gottesman, 1998; Nielsen and Chuang, 2010).
@@
 2.2. The Stabilizer formalism
-A stabilizer of a quantum state $\lvert\psi\rangle$ is an operator $\hat{A}$ such that
-(2) $\hat{A}\ket{\psi}=+1\ket{\psi}$
-Although stabilizers can be chosen from broader operator sets, Pauli stabilizers (i.e., tensor products of the single-qubit operators I, X, Y, Z) are especially popular due to their simple algebraic structure.
-A stabilizer group is a commuting (Abelian) subgroup of the $n$-qubit Pauli group whose elements all square to the identity and do not include the negative identity operator $-I$.
-A state stabilized by a stabilizer group is a stabilizer state.
-Stabilizer states can be equivalently described by the stabilizers in the stabilizer group, especially the group generators.
+[…]
 The effect of Clifford group operations on stabilizer states can be simulated by tracking the transformation of the corresponding stabilizer group.
-Pauli measurement can also be tracked using the following rules (Litinski, 2019),
+Pauli measurement can also be tracked using the following rules (Litinski, 2019):
   1. (1)
@@
   2. (2)
-    if the operator is not in the stabilizer group $\mathcal{S}$, the measurement outcome can be either ±$\pm 1$.
-The state after the measurement is stabilized by a new group that contains the new measured operator ±$\pm g$ depending on the measurement outcome and all other group generators that commute with $g$ in the original stabilizer group $\mathcal{S}$.
-A stabilizer $\mathcal{S}$ represents one possible tensor product of Pauli operators that stabilize a state by equation 2.
+    If the operator is not in the stabilizer group $\mathcal{S}$, the measurement outcome can be either $\pm 1$.
+The state after the measurement is stabilized by a new group that contains the new measured operator $\pm g$ depending on the measurement outcome and all other group generators that commute with $g$ in the original stabilizer group $\mathcal{S}$.
+A stabilizer $\mathcal{S}$ represents one possible tensor product of Pauli operators that stabilize a state.
 However, a pure $n$-qubit stabilizer state is stabilized by $n$ commuting and independent Pauli stabilizers.
@@
 Pauli $X$ is represented by $(1,0)$, $Z\mapsto(0,1)$, $Y\mapsto(1,1)$, and $I\mapsto(0,0)$.
-The commutation relation of two Pauli operators represented as $(x_{1},z_{1})$ and $(x_{2},z_{2})$ can also be easily computed using
-(4) $(x_{1}\cdot z_{2})\oplus(x_{2}\cdot z_{1}),$
-where $\oplus$ is an xor, and a result of 0 means the two Pauli operators commute.
+The commutation relation of two Pauli operators represented as $(x_{1},z_{1})$ and $(x_{2},z_{2})$ can also be easily computed using $(x_{1}\cdot z_{2})\oplus(x_{2}\cdot z_{1})$, where $\oplus$ is an xor, and a result of 0 means the two Pauli operators commute.
 The key purpose of stabilizers in the context of QEC is that errors can easily knock the state out of the stabilizer subspace of the Hilbert space.
@@
 2.3. The CHP Formalism
-Aaronson and Gottesman proposed an optimized version of the tableau, specifically for optimizing the measurement operations, known as the CNOT-Hadamard-Phase (CHP) formalism (Aaronson and Gottesman, 2004).
-For a $n$-qubit state, instead of a single tableau that only tracks the $n$ stabilizer group generators, in CHP formalism, an extra tableau is to added to track the $n$ destabilizers.
+Aaronson and Gottesman proposed an optimized version of the stabilizer tableau for optimizing measurements, known as the CNOT-Hadamard-Phase (CHP) formalism (Aaronson and Gottesman, 2004).
+For an $n$-qubit state, instead of a single tableau that only tracks the $n$ stabilizer group generators, in CHP formalism, an extra tableau is added to track the $n$ destabilizers.
 Specifically, the tableau that are tracked in the CHP formalism can be represented as
-(5) $\left(\begin{array}[]{ccc|ccc|c}x_{11}&\cdots&x_{1n}&z_{11}&\cdots&z_{1n}&r_{1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}&z_{n1}&\cdots&z_{nn}&r_{n}\\ \hline\cr x_{(n+1)1}&\cdots&x_{(n+1)n}&z_{(n+1)1}&\cdots&z_{(n+1)n}&r_{n+1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{(2n)1}&\cdots&x_{(2n)n}&z_{(2n)1}&\cdots&z_{(2n)n}&r_{2n}\end{array}\right)$
+$\left(\begin{array}[]{ccc|ccc|c}x_{11}&\cdots&x_{1n}&z_{11}&\cdots&z_{1n}&r_{1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}&z_{n1}&\cdots&z_{nn}&r_{n}\\ \hline\cr x_{(n+1)1}&\cdots&x_{(n+1)n}&z_{(n+1)1}&\cdots&z_{(n+1)n}&r_{n+1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{(2n)1}&\cdots&x_{(2n)n}&z_{(2n)1}&\cdots&z_{(2n)n}&r_{2n}\end{array}\right)$
 where each row represents a stabilizer or destabilizer.
-The first $n$ rows correspond to the stabilizers, while the last $n$ rows are the destabilizers.
-The columns $r$ represent the sign bits of each Pauli operator.
-By tracking the destabilizers, it removes the need to do Gaussian elimination on the stabilizer tableau when a Pauli measurement is performed, reducing computational complexity from $O(n^{3})$ to $O(n^{2})$ (Aaronson and Gottesman, 2004).
-The tableau requires $4n^{2}+n$ bits to store all necessary stabilizers and destabilizers, including a phase bit for each, for CHP operations.
+The first $n$ rows correspond to the destabilizers, which eliminate the need for $O(n^{3})$ Gaussian elimination (Aaronson and Gottesman, 2004), while the last $n$ rows are the stabilizers.
+The columns $r$ represent the sign bits of each Pauli string (stabilizer).
+The tableau requires $4n^{2}+2n$ bits to store all necessary stabilizers and destabilizers, including a phase bit for each, for CHP operations.
 Thankfully, this means that a full tableau representation of systems with tens of thousands of qubits still only requires space on the order of MB.
-For comparison, a state vector or density matrix representation of the state will grow on the order $O(2^{n})$.
-These simulation methods become infeasible around roughly 50 qubits (Chen et al., 2025).
+For comparison, a state vector or density matrix representation of the state will grow on the order $O(2^{n})$ and $O(4^{n})$ respectively, becoming infeasible with a few dozen qubits (Chen et al., 2025).
 2.4. Operations on the CHP Tableau
@@
 For example, $HXH^{\dagger}=Z$.
-From this, some basic bit transforms on the tableau can be calculated that represent the action of Clifford gates on the stabilizers of a system.
+Some basic bit transforms on the tableau can be calculated that represent the action of Clifford gates on the stabilizers of a system.
 Most often, all stabilizers in the tableau will receive the same transformation gate-by-gate, meaning Clifford gate operations can be similar to matrix operations.
-However, the most time consuming operations, measurements, require more logic.
-Gate Binary Tableau Operation
+However, the most time-consuming operations, i.e., measurements and resets, require much more expensive logic.
+Gate Binary Tableau Operation (Aaronson and Gottesman, 2004)
 Hadamard ($H$) Swap $x$ and $z$; $r=r\oplus(x\cdot z)$
@@
 $r=r\oplus x_{c}\cdot z_{t}$
-Measure ($M$) $\begin{cases}rowsum(i,p)\{\forall i\mid x_{i,q}=1\}&\text{if }random\\ rowsum(temp,i){\forall i>n}&\text{if }deterministic\\ \end{cases}$
+[…]
 Table 1.
@@
 c and t represent the control and target bit of a CNOT operation.
-Other Clifford gates are derived from this basic gate set and minimized.
-Measurement operations in CHP are more complex.
+The original rowsum algorithm is detailed in (Aaronson and Gottesman, 2004) and also provided in Appendix A.2 for convenience.
+Other Clifford gates can be derived from this basic gate set and minimized.
 Performing measurements is sped up by tracking a corresponding destabilizer tableau, but each measurement still requires two distinct steps that branch depending on if the measurement is random or deterministic.
-Algorithm 1 rowsum(h, i) (Aaronson and Gottesman, 2004)
+[…]
+If all stabilizers commute, the measurement outcome is deterministic.
+In this case, all stabilizers stay the same, but a temporary bit array of size $n+1$ is used to collect all of the stabilizer updates.
+rowsum(temp,i) $\forall(i>n)$ collects the outcome from every stabilizer in the tableau.
+Once complete, the phase bit at temp is the measurement outcome.
+2.5. Stabilizer Simulator Execution
+Stabilizer circuits can be split between single-qubit Clifford gates, entangling gates, and measurement gates.
+The time order of all of these gates must be maintained.
+For example, phase and Hadamard gates do not commute, entangling gates need to exchange information at a specific point in time, and measurement gates need all relevant stabilizers to be caught up.
+Time order is more sensitive in the GPU implementation.
+If multiple threads read and write back to the same location in data, changes from one thread often will not be written back before another thread accesses the data.
+Careful management of this aspect was required for Section 3 to provide a speedup without causing inaccuracy due to race conditions.
+Figure.
+2 provides a high-level overview of how circuits can be safely processed on GPU while maximizing efficiency.
+Clifford+M CircuitSingle Gate(q)Measurement(q)Entangling Gate(c,t)SynchronizeExecute CliffordWarp ReductionOutcome DecisionCircuit done?Simulated Quantum StateMCYesNo
+Figure 2. Race condition-safe processing of a Clifford circuit into an binary tableau on GPU.
+2.6. QEC Operations
+One of the most important applications using stabilizer simulation is to simulate QEC operations.
+Specifically, in the syndrome extraction process, the only operations are Clifford gates between syndrome qubits and data qubits, and Pauli-based measurements on those syndrome qubits.
+When we inject Pauli errors with some probability as a model of actual noise on the physical qubits, the process can be simulated completely in stabilizer formalism.
+Over some number of shots to a desired accuracy, these noise injections can simulate or approximate different types of physical error.
+(a) A distance 3 surface code layout
+(b) $X$-checks
+(c) $Z$-checks
+Figure 3. Planar surface code.
+One thoroughly explored QEC code is the planar surface code.
+Planar surface code is a topological QEC code realized on a two-dimensional grid of physical qubits with open boundary conditions, enabling local stabilizer measurements and scalable fault tolerance.
+The qubit layout in a distance $5$ surface code patch is shown in Fig.
+3(a), where the data qubits (blue) are the qubits encoding the quantum information, while $X$ ($Z$) syndrome qubits are to extract the syndrome information for quantum error correction.
+For example, an $X$ ($Z$) syndrome qubit is to measure the stabilizer $\otimes_{j\in\mathcal{N}}X_{j}$ ($\otimes_{j\in\mathcal{N}}Z_{j}$), where $\mathcal{N}$ is the set of neighbor qubits.
+The syndrome extract circuits for X and Z stabilizers are shown in Fig.
+3(b) and 3(c), respectively.
+In a QEC cycle, the X and Z stabilizers of the code patch are extracted, which will be stored and sent to decoders for error correction (Dennis et al., 2002; Fowler et al., 2012a, d, c; Higgott and Gidney, 2025).
+2.7. Stabilizer Noise Modeling
+[…]
+An exact way of simulating a non-Pauli error channel is to decompose it into stochastic Clifford+Reset gates.
+It is not commonly used in QEC investigations for two primary reasons; a) the Pauli frame sampling method, particularly leveraged by Stim for fast noise sampling, is incompatible with stochastic non-Pauli noise, and b) the sampling cost is well known to scale exponentially with the system size and negativity of the distribution (Bennink et al., 2017).
+2.8. Pauli-Based Computing Transpilation
+In recent developments of QEC, especially in superconducting qubit based surface code constructions, the limited connectivity between logical patches has led to the proposal of implementing CX gates via lattice surgery.
+This approach enables straightforward multi-qubit Pauli measurements at the logical level and supports T-gate implementation through multi-patch Pauli measurements combined with magic states.
+Pauli-based computing (PBC) is a quantum computational scheme where qubits undergo a sequence of nondestructive eigenvalue measurements on Hermitian Pauli operators—which may be chosen adaptively based on prior measurement outcomes—and the final result is a classical bit derived from polynomial-time processing of the recorded eigenvalues (Bravyi et al., 2016).
+PBC can readily describe the lattice-surgery-based surface code computation model.
+Specifically, in (Litinski, 2019), a method is introduced to remove all Clifford operations in a quantum circuit by (1) commuting the Clifford gates through the T gates and expanding the T gates into multi-qubit Pauli rotations and (2) absorbing the Clifford operations into measurements to form multi-qubit Pauli measurements.
+The transformation rule can be derived as
+(2) $\left\{\begin{array}[]{l l}P\rightarrow P,&\text{if }[P,P_{c}]=0\\ P\rightarrow iP_{c}P,&\text{if }[P,P_{c}]\neq 0\\ \end{array}\right$.,
+where $P$ is the Pauli-$\pi/4$ rotation axis (represented as a Pauli operator) or the measurement basis, and $P_{c}$ is the Pauli operator corresponding to the Clifford rotation angles (Litinski, 2019).
+2.9. Pauli Grouping
+In quantum algorithms, extracting information from the quantum states prepared by a circuit is crucial, especially in applications such as quantum chemistry, where variational quantum eigensolvers require measuring the expectation value of a Hamiltonian (Peruzzo et al., 2014; Tilly et al., 2022; Cerezo et al., 2021).
+However, quantum hardware natively supports only Pauli-$Z$ measurements on individual qubits.
+Consequently, evaluating the expectation value of problem Hamiltonians requires decomposition into Pauli strings that can be measured with existing hardware.
+To minimize the number of circuit executions shots, a common strategy involves grouping commuting Pauli strings together to measure them simultaneously (Yen et al., 2023).
+Two popular approaches are Qubit-wise commutation (QWC) (Verteletskyi et al., 2020) and Group-wise commutation (GC) (Crawford et al., 2021).
+Suppose we have two $n$-qubit Pauli strings $P_{1}=\bigotimes_{j=1}^{n}\sigma_{k_{j}}^{(j)}$ and $P_{2}=\bigotimes_{j=1}^{n}\tau_{l_{j}}^{(j)}$, where $j$ is the index for qubit, $\sigma_{k_{j}}^{(j)},\tau_{l_{j}}^{(j)}\in\{I,X,Y,Z\}$.
+Under QWC, any two Pauli strings in the same set must satisfy $\forall j$, $\sigma_{k_{j}}^{(j)},\tau_{l_{j}}^{(j)}]=0$, whereas GC only requires $P_{1},P_{2}]=0$.
+Notably, Pauli strings grouped under QWC can be diagonalized using a single layer of single-qubit gates, while GC-based groupings typically require solving for a more complex Clifford circuit to diagonalize the entire set at once.
+3. STABSim Design
+We provide two different implementations of the stabilizer simulator, a lightweight CPU version and a parallelized CUDA version.
+The CPU version implements the tableau as a 2-D grid of independent boolean values for simplicity and quickness.
+Single boolean values in a 2-D array are intuitive to access and manipulate when extending the simulator for tasks beyond direct simulation.
+CHP operations are performed sequentially, bit-by-bit on the tableau using loops.
+While straightforward and fast for small simulations, this approach grows significantly at larger qubit sizes.
+Every Clifford gate update requires a $\theta(n)$ sweep over the tableau to keep the stabilizers up to date, and measurements can be up to $O(n^{2})$ in the worst case.
+For that reason, we focus on STABSim-GPU for our results.
+3.1. The GPU Implementation
+For the GPU implementation, the stabilizer tableau built in STABSim is stored in global memory and accessed by indexing flattened $x$ and $z$ arrays.
+Shared memory is avoided for this task since multi-qubit gates and measurements would require reloading and syncing shared memory across blocks.
+In testing, handling memory between shared blocks meaningfully slowed down circuits with multi-qubit and measurement gates.
+Instead, each thread is responsible for its own stabilizer row of the tableau for non-measurement operations.
+Parallelism across rows avoids race conditions for non-measurement gates and allows the tableau to scale quickly for n-qubits without additional per-thread overhead.
+In contrast to a boolean or packed-bit tableau, which must iterate over a fixed rows for each gate, STABSim’s GPU design parallelizes row updates, keeping latency low as circuit width increases.
+Because a standard CHP designed tableau contains $2n$ rows for an $n$-qubit system, the time to simulate a single Clifford gate increases linearly with qubit count in a CPU simulation.
+Since stabilizers evolve independently under single-qubit Clifford operations, GPU threads can apply gate transformations to each row concurrently, as long as measurement dependencies are accounted for.
+Measurement operations cost the majority of simulation time in dynamic QEC circuits where syndrome extraction occurs in many repeated rounds, in agreement with previous works (Gidney, 2021).
+With that in mind, some key design choices were made to make the most effective use of GPU multi-threading and warp-level primitives.
+3.1.1. Threading
+Threads are responsible for the stabilizers (rows) of the tableau and not the qubits (columns).
+Stabilizers are generally independent of each other in the tableau instruction set until measurement.
+In other words, for bulk Clifford gates, each thread accesses memory in the same pattern but unique locations and performs the same bit operations as every other thread.
+This serves to eliminate the need for synchronization in terms of time-order correctness or any chances of race conditions when rewriting to the tableau.
+3.1.2. Determining the Measurement
+Measurements first require the tableau to be up-to-date to search for a stabilizer that does not commute with the measurement, and that all threads agree on.
+Then measurements need to update potentially the entire remaining $n^{2}$ tableau.
+A few techniques are employed to mitigate the time spent within measurement gates which require costly synchronization bottlenecks.
+The search for a stabilizer that does not commute with the measurement starts with a two step block and grid reduction.
+All blocks find their candidate smallest stabilizer index where x=1 at the gate qubit, for a Z measurement, using $atomicMin$.
+Only blocks that have found a candidate that is less than the number of rows then submit their candidate to a global reduction.
+Using atomic operations here is quite cheap; in a typical QEC construction, a thread will only participate in the atomic operations when an error has occurred and the syndrome measurement is no longer deterministic.
+After those two reductions, a lowest index has been agreed on, with only participation from threads who find anti-commuting stabilizers.
+3.1.3. Warp Primitive Tableau Reduction
+Whether an anti-commuting stabilizer has been found determines the type of measurement outcome.
+In either case, an inner sum of each stabilizer may need to be done followed by a second reduction to a single value if the measurement is deterministic 6.
+To prevent memory-safe but costly atomic operations from queuing and creating a bottleneck, CUDA warp primitives are heavily relied on.
+The random measurement begins with a stride over the number of participating warps.
+This stride is treated as the rows of tableau, since we know every column will participate if a stabilizer row is apart of the measurement update.
+The second step is to filter out non-participant rows, then stride over the number of lanes in a warp, for each column in each row.
+Each lane calculates a polynomial which will be a portion of the power to which $i$ is raised when the Pauli operators represented by the first and second row in a rowsumrowsum iteration are multiplied.
+Each lane can also perform its own updates to the tableau in the random case after this polynomial is calculated.
+This phase is then quickly merged across all warps acting on different columns via __shfl_down_sync\_\_shfl\_down\_sync.
+Random outcomes are simpler upon inspection, because updates happen on independent stabilizer so separate threads can calculate the phase outcome on each row of the tableau.
+Deterministic outcomes are more complicated, since the measurement outcome is calculated recursively as rows are iterated, with dependency on the phase outcome of the last row.
+Furthermore, one scratch row is updated by every participating row in the deterministic measurement, so any way to parallelize this single-index update by multiple rows causes a race condition.
+Deterministic measurements sweep through the tableau to find the final outcome.
+This sweep has to keep track of negative, positive, and zero powers as the phase is modified during calculation, even though the CHP formalism guarantees a single-bit outcome phase at the end.
+To do this, large warp reductions are performed, summing all of the columns of each row, row by row.
+Each block in STABSim contains the maximum number of threads allowable, 1024, to increase access to the same shared memory and maximize where cheaper in-block syncs can be performed, rather than grid-wise syncs.
+Another consideration is that to perform grid-wise syncs without aborting the GPU kernel, the cooperative thread groups primitives allow only one thread block per streaming-multiprocessor (SM).
+Therefore, using the maximum allowable threads offers the best occupancy.
+As 1024 being divisible exactly twice by the number of threads in a each warp, only two warp primitives are needed to rapidly reduce the sum of 1024 unique polynomials, calculated by each thread, down to only one sum per block.
+Since this task stays in-block, its natural to utilize shared memory arrays.
+Finally, the first thread in each block atomically sums with every other block.
+Most often, blocks will have a 0 internal sum and thus will not participate in this atomic sum.
+Still, this atomic will only have nn/1024 participants in the absolute worst case.
+Pseudocode 2 is provided for the two-stage warp reduction to illustrate how the tableau is parsed to maximize the efficiency provided by CUDA warp-level parallelism.
+Expanding rowsumrowsum to a single polynomial helps eliminate thread divergence, aside from functions where only warp or block leaders participate.
+Algorithm 1 Phase Polynomial
+1:function ff($x_{p},z_{p},x_{h},z_{h}$)
+2:  return $x_{p}\cdot\big{(}z_{p}\cdot(z_{h}-x_{h})+(1-z_{p})\cdot z_{h}\cdot(2x_{h}-1)\big{)}$
+3:        $(1-x_{p})\cdot z_{p}\cdot x_{h}\cdot(1-2z_{h})$
+4:end function
+Algorithm 2 Deterministic-$M$ Warp Reduction
+[…]
+2:— Stabilizer Reduction —
+[…]
+4:if anticommutesanticommutes then
+5:  $row\_idx\leftarrow row\cdot cols+i$
+6:  $thread\_sum\leftarrow 0$
+7:  if $i<cols$ then
+8:— Threaded Sum —
+9:    $x_{r}\leftarrow x\_arr[row]$
+10:    $z_{r}\leftarrow z\_arr[row]$
+11:    $x_{s}\leftarrow x\_arr[scratch]$
+12:    $z_{s}\leftarrow z\_arr[scratch]$
+13:    $thread\_sum\leftarrow f(x_{p},z_{p},x_{h},z_{h})$(Alg 1)
+14:    $x\_arr[scratch]\leftarrow x_{s}\oplus x_{p}$
+15:    $z\_arr[scratch]\leftarrow z_{s}\oplus z_{p}$
+16:— Stage 1: Intra-Warp Reduction —
+17:    $warp\_sum\leftarrow\texttt{reduce\_add\_sync}(thread\_sum)$
+18:    if block_lane_id=0block\_lane\_id=0 then
+19:     $\texttt{shared\_mem}[warp\_id]\leftarrow warp\_sum$
+20:    end if
+21:  end if
+22:  syncthreads()
+23:— Stage 2: Intra-Block Reduction —
+24:  if $threadIdx.x<32$ then
+25:    $block\_sum\leftarrow\texttt{reduce\_add\_sync}($
+26:   $\texttt{shared\_mem}[threadIdx.x])$
+27:— Global Reduction —
+28:    if $threadIdx.x=0\land block\_sum\neq 0$ then
+29:     $\texttt{atomicAdd}(global\_sum,block\_sum)$
+30:    end if
+31:  end if
+32:  grid.sync()
+33:  if $i=0$ then
+34:    $r\_arr[scratch\_row]\leftarrow$
+35:   $\begin{cases}0&\text{if }(global\_sum+2r[row]+2r[scratch])\bmod 4=0\\ 1&\text{otherwise}\end{cases}$
+36:    $global\_sum\leftarrow 0$
+37:  end if
+38:  grid.sync()
+39:end if
+3.1.4. Scheduling Experiment
+Parallelizing column wise, i.e.
+assigning other threads to other gates and qubits, effectively shrinks the simulation depth of the circuit, providing another dimension of performance improvement.
+When parallelizing gate operations, gates should be independent of each other to avoid overwriting tableau data while being used by another thread.
+Since individual gate operations are so cheap, usually a handful of xorxor operations, it is easy for threads to jump ahead without synchronization, especially after thousands or millions of operations.
+This idea lends itself to a scheduler, where multiple independent gates are performed simultaneously.
+Scheduling at run time seems useful, but did not benefit our experiments.
+Even the simplest comparisons used to ensure the next gate could safely proceed take longer than the time to directly compute the current gate in the case of a single-pass simulation.
+By defining flags in the circuit at circuit generation time, it could however be preprocessed into safely independent chunks.
+With STABSim-GPU, the user can call $sim(circuit)$ to perform a gate-wise simulation, or sim2d(cirucit,chunk_size)sim2d(cirucit,chunk\_size) to parallelize gates within portions of the circuit where operations are independent from one another.
+In testing QEC-like circuits, little to no advantage was found using gate-parallelization.
+QEC runtime is dominated by the time to evaluated mid-circuit measurements, so improving the extremely cheap chunks of non-measurement gates in between had little to no effect.
+This function is still included in STABSim because better suited workloads could utilize parallelized Clifford gates not separated by measurement to speed up simulation time (Osama et al., 2025).
+3.2. Pauli Grouping Using STABSim
+Pauli grouping is a well used and understood method in quantum chemistry and optimization.
+Because of this usefulness, and applications to algorithms 4 and 5 in our T-Gate transpiler, we quickly discuss functions for grouping in the context of binarized stabilizers.
+Using $x$ and $z$ as the bits of the $X$ and $Z$ components of the stabilizers (Pauli strings), $(x^{i}_{1}\land z^{i}_{2})\oplus(x^{i}_{2}\land z^{i}_{1})=0$ if they commute at each qubit $i$.
+Running an xor product can extend the commutation check from qubit-wise to group-wise over the full Pauli string:
+(3) $\bigoplus_{i=1}^{n}((x^{i}_{1}\land z^{i}_{2})\oplus(x^{i}_{2}\land z^{i}_{1}))=0$
+if $P_{1}$ and $P_{2}$ commute.
+If the binary vectors can be packed into larger data types (e.g.
+32, 64-bit words), they can be quickly processed with vectorized bitwise instructions.
+For group-wise commutation, the xor product reduction can be replaced in favor of a population count on the vectorized results if its an available intrinsic.
+For example, if
+(4) $\mathrm{popcount}\left(\left\{i\in[n]\;\middle|\;[P_{1}^{(i)},P_{2}^{(i)}]\right\}\right)\bmod 2=0,$
+then $P_{1}$ and $P_{2}$ group-commute.
+To evaluate the commutation of a Pauli string $P$ against an entire stabilizer tableau using matrix operations, the following matrix expression for all stabilizers $S$ in the tableau is used:
+(5) $(x_{P}\cdot Z_{S})+(z_{P}\cdot X_{S})\mod 2=0$
+if $S$ commutes with the full tableau.
+Here, $x_{P}$ and $z_{P}$ are the binary row vectors representing the $X$ and $Z$ components of the Pauli string $P$.
+$X_{S}$ and $Z_{S}$ are the binary matrices corresponding to the $X$ and $Z$ components of all stabilizers $S$ in the tableau.
+The final sum matrix is an $n\times n$ matrix of the qubit-wise commutation results from $P,S_{i}]$ where $i$ is each row.
+3.3. Clifford Push Through Transpiler
+Transpiling a Clifford+T universal quantum circuit into Pauli measurements with a separate T-layers has great potential to make efficient use of quantum machinery by abstracting away Clifford operations, and minimizing expensive magic state operations.
+Commuting T-layers can be optimized after separation by removing redundant eighth rotations and pushing them to the Clifford layer as quarter rotation gates.
+3.3.1. Measurement Tableau and T-Tableau Construction
+With the target Clifford+T circuit arranged in time order, gates are extracted from the end and added to one of two sets of tableaus.
+The first tableau is an n-qubit identity tableau of only stabilizers, used for tracking Clifford operations absorbed into the measurement.
+The second is an empty n-qubit tableau used for tracking the commutation rules of T gates as they are pushed through the circuit.
+If the gate extracted is a T gate, it is appended to the end of the T tableau as an n-qubit Pauli string with a Z stabilizer at the qubit index of the T gate.
+If the gate extracted is a Clifford operation, it is applied to the measurement (M) tableau by CHP gate operations.
+It is also applied to the T tableau by the same rules, but only acts on the T gates which have been appended up to that point.
+For example, if from the end of the original circuit, no T gates have been extracted yet, then the Clifford gate will have no affect on the T tableau as there are no interactions with the T gates that come before it in time order.
+Detailed tableau construction psuedocode is provided in Appendix 3.
+3.3.2. T-Layer Separation
+Once the full circuit has been parsed and M and T tableaus have been created, the T tableau is split into commuting groups as follows.
+Starting from the end of the T tableau (the last appended stabilizer, i.e.
+the first T gate in time order) a new empty tableau that we will call $P_{0}$ is made and the stabilizer is appended.
+Each subsequent stabilizer from the end of the T tableau is checked against the new P tableau using the commutation method discussed in  3.2.
+If the subsequent stabilizer commutes with each Pauli string in the first P tableau, it is appended to that tableau.
+If there is an anti-commutation between the stabilizer being checked and a stabilizer within the tableau, a new tableau ($P_{1}$) is created.
+For the next stabilizer from the end of T, it is first checked against $P_{0}$.
+If it commutes with $P_{0}$ it is appended to $P_{0}$.
+If not, the stabilizer is checked against $P_{1}$ and appended to $P_{1}$ if it commutes.
+If it anti-commutes, the stabilizer starts a new tableau $P_{2}$.
+This process continues from the end of T for all stabilizers in T until T is empty.
+After this process is complete, there will be some number of self-commuting P tableaus (at least one, at most the number of T gates in the original circuit).
+Figure 4.
+T-Separation.
+The T-Gate tableau is separated into commuting groups by placing the first T-stabilizer at the end of the original circuit order.
+If the stabilizer does not commute, it cannot be grouped into the same T-layer, and must be appended to an earlier T-Layer or create its own.
+At this point, the basic measurement circuit and T layers have been created.
+The total number of T-stabilizers is counted to later determine when optimization is done.
+Clifford rotations are now extracted from the P tableaus with the goal of reducing the depth of the T layers.
+To do this, repeated stabilizers within each tableau are counted, starting from the last P tableau.
+The P tableaus are in reverse time order, meaning the first entry of $P_{0}$ contains the very last T gate in the original circuit.
+We want to push Clifford gates to the end, towards the measurement circuit, to absorb any leftover Clifford rotation.
+For every two occurrences of a stabilizer, a quarter rotation gate can be extracted.
+Likewise, for every four occurrences of a quarter rotation gate, a full rotation is made, and all eight T stabilizers can be removed without pushing any operations through to the measurement tableau.
+3.3.3. T-Layer Optimization
+We start by rightward gate pushing from the last P tableau, i.e., the first T gates in original circuit order.
+Repeat stabilizers in the tableau are counted, removed from the tableau, and the corresponding quarter rotations are then checked For the second to last P tableau, repeat stabilizers are once again counted and removed, but now need to be commuted through the P tableau to the right before being added to the beginning of the measurement circuit.
+To do this, one copy of the repeated stabilizer is appended to the next P tableau and checked to see if it commutes with every stabilizer in the tableau.
+If it does, it is passed through to the next tableau.
+If not, a rowsumrowsum(algorithm 6) is called on the appended stabilizer with each row of the tableau that does not commute.
+rowsumrowsum accumulates commutation the effect of an anti-commuting Pauli string throughout the stabilizer tableau by summing the contributions of each stabilizer Pauli and corresponding phase bit.
+The sum modulo 4 of the appended Pauli string respective non-commuting stabilizer becomes the new phase bit $r$ of that stabilizer.
+rowsumrowsum is originally used in the CHP formalism to propagate the effect of measurement on the tableau.
+To satisfy the commutation rules for pushing a Clifford rotation gate through another arbitrary rotation (Litinski, 2019), a phase of $i$ is added to the rowsumrowsum product.
+Figure 5. T Optimization. Every two repeating stabilizers are removed from the tableau and pushed left, following the $rowsum+i$ routine for each unique Pauli string.
+Once a pass of the T-optimization algorithm is complete, the total T-stabilizers remaining are counted.
+If no repeat stabilizers were extracted during optimization, the process ends.
+Otherwise, optimization is repeated to extract new repeat stabilizers created by $rowsum+i$, until the total number of T-stabilizers no longer changes.
+After optimization, the $P$ tableaus can be recombined to a larger $T$ tableau.
+The stabilizers of the $T$ tableau are the self-entangled layers of the $\pi/8$ rotation gates that remain.
+In Figure 2, every non-Identity Pauli operator in all the layers is counted as a T-rotation, for comparison with the number states needed for T-gates in the original circuit.
+All of our protocols detailed in Section 3.3 are also rewritten in psuedo-code in Appendix A.1 for further reference.
+4. Evaluation
+CPU simulator evaluations were all performed on an AMD EPYC 7502 compute node with a 3.35 GHz boost clock and 256 GB of RAM.
+GPU evaluations were performed on an NVIDIA H100 SXM node.
+These platforms were chosen for a stable and consistent test environment versus a laptop or desktop that may have power saving modes and background processes.
+STABSim-GPU used CUDA 12.1.
+4.1. QEC Simulation
+Stabilizer simulation is made feasible in time and size by the relatively slow growth of the stabilizer tableau as qubit count increases compared to other simulation methods.
+However, measurement operations are $O(n^{2})$ worst-case in the original CHP formalism, one dimension due to every stabilizer potentially being checked, and another from every column participating in rowsumrowsum when the row check conditions are met.
+Measurements greatly bottleneck the time for dynamic circuits, especially QEC codes where many repeated stabilizer measurements are the primary interest for evaluation.
+Figure 6.
+Full tableau simulation time of the planar surface code as code distance increases.
+Procedural generation highlights faster per-gate simulation times as opposed to deeper targeted compilation.
+The circuit here follows the example given in Section 2.6.
+Stim improves on the deterministic measurement worst case, and, combined with 256-bit SIMD instructions, we find that STABSim starts outperforming Stim at distance 11 in the planar code.
+Beyond distance 11, STABSim scales much better than existing simulators as the cost per qubit is almost constant.
+STABSim’s scaling is most strongly a function of the number of measurement gates.
+The primary cost is in synchronization to prepare for warp reduction once all threads have done their computation, which is still quite cheap compared to a $O(n)$ or $O(n^{2})$ overhead per gate.
+Figure 7 shows how different simulators scale in a wide circuit random benchmark, where a random gate selected out of H, S, CX, Measurement, and Reset are applied in a number of rounds equal to the number of qubits.
+This benchmark highlights the polynomial cost of the stabilizer tableau.
+STABSim gets much closer to a linear per-gate cost, versus polynomial gate times in a sequential simulation.
+Stim and Qiskit both compiled the circuit automatically before simulation, causing deviations depending on how much they were able to simplify each circuit.
+STABSim ran the original circuit directly.
+Figure 7. Full tableau simulation of a randomly generated $n$ width circuit with $n$ rounds. This figure highlights performance in scaling with more qubits.
+4.2. Better Noise Modeling
+As discussed in 2.7, existing noise modeling methods generally must either make nonphysical approximations to Pauli channels, or suffer an exponential overhead that grows with qubits (through quasi-probability sampling or density matrix simulation).
+However, by compositing the qubit relaxation channel with the qubit phase decoherence channel into a single quasi-distribution,
+$\mathcal{E}_{phase}(\mathcal{E}_{relaxation}(\rho)),$
+we implement new a $T1$ and $T2$ noise model that has both constant factor overhead, like SPEC, but also exact accuracy when $T1\geq T2$.
+Figure 8 shows an analytical evaluation of how the known quasi-probability decomposition for relaxation and decoherence, which can be derived from (Bennink et al., 2017), also implemented in STABSim, compares to our composite distribution.
+When $T1\geq T2$, the composite probability is entirely positive, and can be sampled cheaply with Monte Carlo methods.
+The number of shots required beyond a completely positive baseline distribution is $\propto\Gamma^{2}$, where $\Gamma$ is the overhead factor of the negative quasi-distribution, defined as $1+2\mathcal{N}$, or $\sum_{p}{|p|}$ for all probabilities in the quasi-distribution $\mathbb{P}$.
+Figure 8.
+Samples to acheive equal precision, $n$ = qubits.
+Left: Selected $\mathbf{\tau}$ show how sampling overhead is proportional to the negativity $\mathcal{N}$ which is a function of the ratio of $\mathbf{\tau}$, the error duration, to $\mathbf{T1}$, the relaxation time.
+The composite probability distribution $\mathbb{P}$ is positive when $T1\geq T2$, meaning it can be sampled without variance overhead.
+Right: Sweeping over $\mathbf{\tau}$ in relation to $T1$ visualizes how sampling cost scales with $\mathbf{\tau}$ with prior methods when $T1\geq T2$.
+4.3. Pauli Grouping
+To demonstrate a chemistry utility of STABSim, we apply STABSim to Pauli grouping problems.
+We focus on grouping the Pauli strings from molecular Hamiltonians, which are widely used in quantum chemistry applications.
+Hamiltonians were chosen for varying degrees of freedom and measurement gate sizes, while remaining practical to simulate with our state vector simulator.
+The detailed information about the molecular configuration settings and the Pauli group terms are shown in Table 2.
+To group the Pauli strings, the following steps were employed.
+First, Pauli terms are sorted up front by the weight of the Pauli term coefficients.
+Groups of commuting Pauli terms are then greedily constructed starting with the highest weight terms by either QWC or GC.
+This results in higher weight terms being grouped earlier so sampling resources can be prioritized to get a higher fidelity from more shots on larger terms.
+Table 2 counts the total number of grouped gates.
+To verify the correctness of the ground state energy result after grouping, all molecular Hamiltonians were evaluated with a statevector simulator we have developed.
+Our tableau-based simulator enables fast and scalable implementation of both strategies with xor product commutation checks.
+The time spent performing grouping and generating measurement settings was negligible compared to the overall simulation time, especially when considering the time for statevector-based energy verification.
+VQE in particular still faces challenges in exponential post-processing complexity (Scriva et al., 2024), but Pauli grouping is a useful and scalable pre-processing step that can be used with any Pauli-term based quantum algorithm to reduce resource overhead.
+Molecule $n$ Initial QWC GWC % Change
+H4 4 3428 717 275 61.65%
+LiH 3 756 449 316 29.62%
+LiH 4 3428 878 427 51.37%
+H6 18 16196 3844 1858 51.66%
+H6 22 16196 3850 1946 49.45%
+BeH2 6 16196 1681 749 55.44%
+BeH2 7 27124 3627 1580 56.44%
+Table 2.
+Gates required for evaluation after grouping small molecular Hamiltonians.
+Initial basis gates, Qubit-Wise, and Group-Wise gates are shown, along with the % improvement from QWC to GWC in evaluation gate count.
+$n$ is the number of qubits/orbitals considered in VQE.
+4.4. PBC Transpilation
+Like Pauli grouping, the T-Transpiler application relies on fast bitwise commutation checks to be very efficient.
+Figure 9 shows the total time STABSim took to perform Algorithms 34, and 5.
+The purpose of the T-transpiler is to allow for easier scheduling of T-gate factories in quantum architectures by a) lowering the number of T layers, and b) providing an equal cost for each layer.
+PyZX contains the same functionality, but converts the circuit to a ZX-calculus representation that can then be minimized (Kissinger and Wetering, 2020).
+For all of the circuits tested, both STABSim and PyZX converged on the same T-optimization factor.
+The difference in time is how long it took to arrive at that factor.
+Figure 9. Lower is faster. Transpilation $log_{10}$ time for STABSim using the design discussed in Section 3.3 vs PyZX optimizing on ZX-calculus (Wetering, 2020). Both converge on the same T-counts.
+Most of the substantial time gain found with STABSim is an improvement in the optimization step.
+Processing is primarily an issue of circuit size, and does not require commutation checks or other applications that can be sped up by bitwise methods.
+A complete table of optimized Clifford+T to PBC circuit statistics is provided in Appendix A.3.
+The T-stabilizers remaining after optimization comprise most of the remaining circuit, apart from the now-merged and rotated measurement operations at the end.
+Each stabilizer in the T-tableau after the full processing is an entangled layer of $\pi/8$ rotational gates in the X, Y, or Z basis, given by the Pauli operator on each qubit.
+A phase bit of 1 in the stabilizer means the basis rotation in that layer is negative.
+The circuits with more trivial T-States tended to converge quickly and eliminate many T-layers, which reduced the size and frequency of loops in Algorithm 5, where commutation checks become a bottleneck.
+5. Related Work
+Many projects have been developed on the concept of simulating a quantum system from its stabilizers since CHP was proposed by Aaronson and Gottesman (Aaronson and Gottesman, 2004).
+Cirq, Qiskit, and other projects often implement stabilizer simulators as functions in their libraries (Developers, 2021; Qiskit Development Team, 2021).
+More recently, the Stim project made large gains in QEC sampling via meticulous CPU-based optimization and Pauli frame simulation alongside the stabilizer tableau (Gidney, 2021).
+Those tools are the most comparable to STABSim, as we aim to scale up and expand the functions they perform.
+Cuda-QEC leverages GPU for accelerated decoding of error syndromes (NVIDIA Corporation & Affiliates, [n.
+d.]).
+However, it does not provide a stabilizer simulator for characterizing large error correcting codes.
+Rather, Cuda-QEC focuses on parallelism for decoding syndromes, which much be provided by a simulator.
+Qua-SARQ performs GPU accelerated simulation of Clifford gates on a stabilizer tableau, to replace the work of Stim in the equivalence checking applications(Osama et al., 2025).
+Crucially, Qua-SARQ only simulates Clifford gates, which, given the similarity to large binary matrix operations, are a portion of stabilizer simulation that are straightforward to parallelize.
+Qua-SARQ does not address measurement gates, which turn out to be the most costly portion of simulation and the crux of syndrome extraction in QEC (Gidney, 2021).
+Pauli measurements and Pauli-Based Computation (PBC) are discussed significantly by Litinski in the context of surface code operations for fault-tolerance (Litinski, 2019).
+Multiple works explore PBC compilation using various methods, including a standalone Python compiler using space/time tradeoffs with auxiliary qubits, and a ZX calculus method implemented in PyZX (Peres and Galvão, 2023; Kissinger and Wetering, 2020).
+Our work is the first to formulate PBC in the context of a stabilizer tableau, where highly efficient bitwise stabilizer operations are exploited for much faster compilation than existing methods.
+6. Conclusion
+Large stabilizer simulations become more and more of interest as quantum computing moves towards intermediate scale and eventual fault tolerance.
+Applications benefiting from large scale stabilizer-based computing can be found in higher-distance error correcting codes, quantum verification, limited Clifford+T circuits, heterogeneous codes, and more (Bravyi et al., 2024b; Fowler et al., 2012b; Bravyi and Gosset, 2016; Stein et al., 2024).
+STABSim shows promise in evaluating growing QEC codes or other large circuits in reasonable amounts of time by scaling better than any existing stabilizer simulator.
+We furthermore show how the extended capabilities of STABSim directly benefit other domains beyond direct simulation.
+Work in these areas has also inspired ideas for future directions.
+For example, application of our Clifford simulator to other non-Clifford channels decomposed into quasi-probability distributions is interesting, as it could explore a new regime of practical simulation via rapid Monte Carlo sampling.
+GPU has the potential to also perform parallel sampling, if memory costs can remain practical.
+Investigation beyond the scope of this work could provide better guidance on which error sampling methods are best to use in different scenarios.
+Acknowledgment
+This research was supported by the Quantum Algorithms and Architecture for Domain Science Initiative (QuAADS), under the Laboratory Directed Research and Development (LDRD) Program at Pacific Northwest National Laboratory (PNNL).
+M.
+Wang and A.
+Li were supported by the U.S.
+Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (QSC).
+This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.
+This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S.
+Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No.
+DE-AC02-05CH11231.
+The Pacific Northwest National Laboratory is operated by Battelle for the U.S.
+Department of Energy under Contract DE-AC05-76RL01830.
+Appendix A Appendix
+This appendix provides detailed pseudocode for the algorithms used in transpilation, the T-optimization data set, and the original CHP rowsumrowsum algorithm for convenience to refer to throughout.
+A.1. Tableau T-Transpiler Algorithms
+The following are the three key stages described in Section 3.3 and transcribed here into pseudocode for further reading.
+Algorithm 3 Tableau Construction
+1:Initialize an empty tableau $T_{tab}$ to store $pi/4$ rotation gates ($T$ and $T^{\dagger}$), and an $n\times n$ identity tableau $M_{tab}$.
+2:for all gates $G$ at qubit $q$ starting from the end do
+3:  if $G=T_{q}$ then
+4:   Append an $n$-qubit $Z_{q}$-stabilizer to $T_{tab}$ and set $r\leftarrow 0$
+5:  else if $G=T^{\dagger}_{q}$ then
+6:   Append an $n$-qubit $Z_{q}$-stabilizer to $T_{tab}$ and set $r\leftarrow 1$
+7:  else if $G$ is Clifford then
+8:   Apply $G$ to $M_{tab}$ and $T_{tab}$ using CHP rules
+9:  end if
+10:end for
+Algorithm 4 T-Separation
+1:Initialize tableau list $\mathcal{P}\leftarrow[\ ]$
+2:Create tableau $P_{0}$ and append the last stabilizer in $T_{\text{tab}}$
+3:Append $P_{0}$ to $\mathcal{P}$
+4:for all stabilizers $S$ from second-to-last to first in $T_{\text{tab}}$ do
+5:  $appended\leftarrow\textbf{false}$
+6:  for all $P\in\mathcal{P}$ starting from $P_{0}$ do
+7:   if $[S,P]=0$ then
+8:     Append $S$ to $P$
+9:     $appended\leftarrow\textbf{true}$
+10:     break
+11:   end if
+12:  end for
+13:  if not appendedappended then
+14:   Create new tableau $P_{\lvert\mathcal{P}\rvert}$ and append $S$
+15:   Append $P_{\lvert\mathcal{P}\rvert}$ to $\mathcal{P}$
+16:  end if
+17:end for
+Algorithm 5 T-Optimization
+Count initial total number of rows in all $P_{i}\in\mathcal{P}$, store as $N_{\text{prev}}$
+Initialize convergence flag $converged\leftarrow\textbf{false}$
+while not convergedconverged do
+  for $i=\lvert\mathcal{P}\rvert-1$ down to 0 do
+   while there exist two identical stabilizers $S$ in $P_{i}$ do
+     Remove both copies of $S$ from $P_{i}$
+     Append one copy of $S$ to $P_{i+1}$
+     for all stabilizers $S_{T}$ in $P_{i+1}$ do
+      if $S_{T},S]\neq 0$ then
+        Apply $\text{rowsum}(T,S)$
+      end if
+     end for
+     Remove $S$ from $P_{i+1}$ and append to $P_{i+2}$
+   end while
+  end for
+  Count the total number of rows in all $P_{i}\in\mathcal{P}$, store as $N_{\text{new}}$
+  if $N_{\text{new}}=N_{\text{prev}}$ then
+   $converged\leftarrow\textbf{true}$
+  else
+   $N_{\text{prev}}\leftarrow N_{\text{new}}$
+  end if
+end while
+$\lvert\mathcal{P}\rvert$ is the number of $\pi/8$ rotation gate layers after minimization.
+The sum of all non-Identity Pauli operators in all $P_{i}$ stabilizers is the number of $\pi/8$ rotation gates.
+A.2. Rowsum Algorithm
+The original CHP rowsumrowsum algorithm is key to understanding the parallel implementation used to determine stabilizer phase.
+This algorithm makes use of a helper function $g$ for computing local commutation contributions.
+Algorithm 6 rowsum(h,i)(h,i)
 1:function g($x_{1},z_{1},x_{2},z_{2}$)
 2:  if $x_{1}=0$ and $z_{1}=0$ then
-3:         return 00
+3:   return 0
 4:  else if $x_{1}=1$ and $z_{1}=1$ then
@@
 27:end function
-The first step is to search for a stabilizer that does not commute with the measurement.
-A non-commuting stabilizer indicates a random measurement outcome.
-In that case, only non-commuting stabilizers are updated.
-p is the first stabilizer that does not commute with the measurement.
-rowsum(i,p)$\{i\mid x_{q,i}=1\}$ updates the phase bit of any non-commuting row i using row p.
-The current stabilizer at p is then stored in the corresponding destabilizer, and the stabilizer at p is projected to the measurement basis, Z in this case.
-Finally a random phase bit for p, i.e.
-the measurement outcome, is chosen.
-If all stabilizers commute, the measurement outcome needs to be determined.
-In this case all stabilizers stay the same, but a temporary bit array of size $n+1$ is used to collect all of the stabilizer updates.
-rowsum(temp,i)$\forall(i>n)$ collects the outcome from every stabilizer in the tableau.
-Once complete, the phase bit at temp is the measurement outcome.
-2.5. Execution of the Stabilizer Simulator
-Stabilizer circuits can be split between single qubit Clifford gates, entangling gates, and measurement gates.
-The time order of all of these gates is critical for correctness.
-For example, phase and Hadamard gates do not commute, entangling gates need to exchange information at a specific point in time, and measurement gates need all relevant stabilizers to be caught up.
-Time order is more sensitive in the GPU implementation.
-If multiple threads read and write back to the same location in data, changes from one thread may not be written back before another thread accesses the data.
-In STAB-GPU, every thread is responsible for its own stabilizer row.
-As such, single qubit gates can be run without synchronization, since there is no risk of another thread updating a given stabilizer before the previous thread finishes an update.
-Entangling gates rely on and swap information in multiple columns, but since each thread is responsible for dedicated rows, there is no conflict until measurement.
-Measurement gates require that the circuit execution is caught up before proceeding.
-Even more costly, execution has to pause while measurement gates perform intermediate calculation that can easily be overwritten by other threads executing later gates.
-Luckily there is parallelism to be exploited in measurements discussed in section 3.1 to offset the cost of grid-wise synchronization.
-Clifford Circuit – GatesSingle Qubit Gate(q)Measurement(q)Entangling Gate(c,t)c, t in the same block?Synchronize BlockSynchronize GridExecute CliffordExecute MeasurementWaitAll gates processed?Updated TableauNoYesM2-Qubit GateYesNo
-Figure 2.
-Gate-by-gate processing of a Clifford circuit into an binary tableau on GPU.
-Single qubit targets are denoted by q, while c and t denote the control and target qubits of a two-qubit Clifford gate.
-2.6. QEC Operations
-One of the most important applications using stabilizer simulation is to simulate QEC operations.
-Specifically, in the syndrome extraction process, the only operations are Clifford gates between syndrome qubits and data qubits and Pauli base measurements on the syndrome qubits.
-When we inject Pauli errors as a model of actual noise on the physical qubits, the process can be simulated completely in stabilizer formalism.
-(a) A distance 3 surface code layout
-(b) $X$-checks
-(c) $Z$-checks
-Figure 3. Planar surface code.
-Here, we take plainer surface code as an example.
-Planar surface code is a topological quantum error-correcting code realized on a two-dimensional grid of physical qubits with open boundary conditions, enabling local stabilizer measurements and scalable fault tolerance.
-The qubit layout in a distance $5$ surface code patch is shown in Fig.
-3(a), where the data qubits (blue) are the qubits encoding the quantum information, while $X$ ($Z$) syndrome qubits are to extract the syndrome information for quantum error correction.
-For example, an $X$ ($Z$) syndrome qubit is to measure the stabilizer $\otimes_{j\in\mathcal{N}}X_{j}$ ($\otimes_{j\in\mathcal{N}}Z_{j}$), where $\mathcal{N}$ is the set of neighbor qubits.
-The syndrome extract circuits for X and Z stabilizers are shown in Fig.
-3(b) and 3(c), respectively.
-In a QEC cycle, the X and Z stabilizers of the code patch are extracted, which will be stored and sent to decoders for error correction (Dennis et al., 2002; Fowler et al., 2012a, d, c; Higgott and Gidney, 2025).
-2.7. Pauli-Based Computing Transpilation
-In the recent development of QEC, especially in superconducting-circuit-based surface code constructions, the limited connectivity between logical patches has led to the proposal of implementing CX gates via lattice surgery.
-This approach enables straightforward multi-qubit Pauli measurements at the logical level and supports T-gate implementation through multi-patch Pauli measurements combined with prepared magic states.
-Pauli-based computing (PBC) is a quantum computational scheme where qubits undergo a sequence of nondestructive eigenvalue measurements on Hermitian Pauli operators—which may be chosen adaptively based on prior measurement outcomes—and the final result is a classical bit derived from polynomial-time processing of the recorded eigenvalues (Bravyi et al., 2016).
-PBC can readily describe the lattice-surgery-based surface code computation model.
-Specifically, in (Litinski, 2019), a method is introduced to remove all Clifford operations in a quantum circuit by (1) commuting the Clifford gates through the T gates and expanding the T gates into multi-qubit Pauli rotations and (2) absorbing the Clifford operations into measurements to form multi-qubit Pauli measurements.
-The transformation rule can be derived as
-(6) $\left\{\begin{array}[]{l l}P\rightarrow P,&\text{if }[P,P_{c}]=0\\ P\rightarrow iP_{c}P,&\text{if }[P,P_{c}]\neq 0\\ \end{array}\right.,$
-where $P$ is the Pauli-$\pi/4$ rotation axis (represented as a Pauli operator) or the measurement basis, and $P_{c}$ is the Pauli operator corresponding to the Clifford rotation angles (Litinski, 2019).
-2.8. Pauli Grouping
-In quantum algorithms, extracting information from the quantum states prepared by a circuit is crucial, especially in applications such as quantum chemistry, where variational quantum eigensolvers require measuring the expectation value of a Hamiltonian (Peruzzo et al., 2014; Tilly et al., 2022; Cerezo et al., 2021).
-However, quantum hardware natively supports only Pauli-$Z$ measurements on individual qubits.
-Consequently, evaluating the expectation value of problem Hamiltonians requires decomposing them into Pauli strings that can be measured with existing hardware.
-To minimize the number of circuit executions shots, a common strategy involves grouping commuting Pauli strings together to measure them simultaneously (Yen et al., 2023).
-Two popular approaches are Qubit-wise commutation (QWC) (Verteletskyi et al., 2020) and Group-wise commutation (GC) (Crawford et al., 2021).
-Suppose we have two $n$-qubit Pauli strings $P_{1}=\bigotimes_{j=1}^{n}\sigma_{k_{j}}^{(j)}$ and $P_{2}=\bigotimes_{j=1}^{n}\tau_{l_{j}}^{(j)}$, where $j$ is the index for qubit, $\sigma_{k_{j}}^{(j)},\tau_{l_{j}}^{(j)}\in\{I,X,Y,Z\}$.
-Under QWC, any two Pauli strings in the same set must satisfy
-(7) $\forall j,\ [\sigma_{k_{j}}^{(j)},\tau_{l_{j}}^{(j)}]=0,$
-whereas GC only requires $[P_{1},P_{2}]=0$.
-Notably, Pauli strings grouped under QWC can be diagonalized using a single layer of single-qubit gates, while GC-based groupings typically require solving for a more complex Clifford circuit to diagonalize the entire set at once.
-3. STABSim Design
-We provide two different implementations of the stabilizer simulator, a lightweight CPU version and a parallelized CUDA version.
-The CPU version implements the tableau as a 2-D grid of independent boolean values for simplicity and quickness.
-Single boolean values in a 2-D array are intuitive to access and manipulate when extending the simulator for tasks beyond direct simulation.
-CHP operations are performed sequentially, bit-by-bit on the tableau using loops.
-While straightforward and fast for small simulations, this approach grows significantly at larger qubit sizes.
-Every Clifford gate update requires a $\theta(n)$ sweep over the tableau to keep the stabilizers up to date, and measurements can be up to $O(n^{2})$ in the worst case.
-For that reason, we focus on STABSim-GPU for our results.
-3.1. GPU implementation
-For GPU implementation of STABSim, the stabilizer tableau built in STAB-Sim is stored in global memory and accessed by indexing flattened $x$ and $z$ arrays.
-Shared memory is avoided for this task since multi-qubit gates and measurements would require reloading and syncing shared memory across blocks.
-Despite faster memory access in theory, handling memory between shared blocks slowed down circuits with meaningful multi-qubit and measurement gates.
-Instead, each thread is responsible for its own stabilizer row of the tableau for non-measurement operations.
-Row-level parallelism avoids race conditions for non-measurement gates and allows the tableau to scale quickly without more operations per thread.
-In contrast to a boolean or packed-bit CPU tableau, which must iterate over rows for each gate, STAB-Sim’s GPU design parallelizes row updates, significantly reducing latency for wide or deep circuits.
-Because a stabilizer tableau contains 2n rows for an n-qubit system, the time to simulate a single gate increases linearly with qubit count in a sequential simulation.
-However, since stabilizers evolve independently under single-qubit Clifford operations, GPU threads can apply gate transformations to each row concurrently as long as multi-qubit gate dependencies are accounted for.
-STAB-GPU also supports $sim2D$, a simulation method which computes all user-defined independent gates at once.
-This is useful to the extent that the target circuit is limited by the depth of Hadamard, phase, and CNOT gates.
-However, performance gains depend heavily on circuit structure.
-For example, applying $sim2D$ to the ancilla initialization and entanglement in a surface code cycle 7 resulted in negligible speedup, likely due to the high frequency of intermediate measurements limiting gate parallelism.
-On the other hand, $sim$ should be used for circuits with dynamic measurement operations.
-Measurement operations particularly comprise the bulk of the simulation time, especially in any circuit where measurements are more frequent than a single end layer.
-In STAB-GPU, measurement operations are scheduled with block and grid-wide synchronization steps and occasionally utilize atomic operations to collaborate.
-To start any measurement, the simulator first searches for stabilizer in the tableau that does not commute with a Z-measurement at the target qubit.
-In practice, this is a search for the first stabilizer row where $x=1$ at the target qubit column.
-Every thread calculates its own row index, and the first row where $x=1$ is determined by a global $atomicMin$ of all threads after an initial inner-block reduction.
-If an anti-commuting stabilizer is found, the measurement will be deterministic, and random if not.
-In both cases, the outcome and tableau update are computed using the $rowsum$ operation introduced by Aaronson and Gottesman (Aaronson and Gottesman, 2004) for CHP, which is parallelized in STAB-GPU.
-When the measurement is deterministic, $rowsum$ aggregates all relevant stabilizers into a single scratch row to calculate the measurement result.
-The number of row operations is worst case $O(n)$ if every stabilizer anti-commutes.
-Deterministic measurement outcomes rely on stabilizer rows to be aggregated in order, and the inner sum is calculated column by column in $rowsum$.
-This portion is $\Theta$(n) since every column must participate in the check, so parallelization inside $rowsum$ has high utilization.
-Each thread calculates its contribution to $rowsum$ based on its unique column bit in the source and target rows.
-Once that is done, every thread contributes to an atomic $rowsum$, where the final $r$ bit is determined by all of the contributions.
-For random outcomes, the direction of $rowsum$ is reversed: the selected stabilizer row propagates its update across anti-commuting rows.
-$rowsum$ is still the dimension being parallelized since its always linear time, but some of the synchronization requirement is removed since the updates to being broadcast to independent stabilizers are independent of each other.
-3.2. Scheduling Experiment and Gate-Wise Parallelization
-Parallelizing column wise, i.e.
-assigning other threads to other gates and qubits, effectively shrinks the simulation depth of the circuit, providing another dimension of performance improvement.
-When parallelizing gate operations, gates should be independent of each other to avoid overwriting tableau data while being used by another thread.
-Since individual gate operations are so cheap, usually a handful of $xor$ operations, it is easy for threads to jump ahead without synchronization, especially after thousands or millions of operations.
-This idea lends itself to a scheduler, where multiple independent gates are performed simultaneously.
-Scheduling at run time seems useful, but did not benefit in our experiments.
-Even the simplest comparisons used to ensure the next gate could safely proceed take longer than the time to directly compute the current gate in the case of single-pass simulation.
-By defining flags in the circuit at circuit generation time, it can be preprocessed into chunks safely independent chunks.
-With $StabGPU$, the user can call $sim(circuit)$ to perform a gate-wise simulation, or $sim2d(cirucit,chunk\_size)$ to parallelize gates within portions of the circuit where operations are independent from one another.
-In testing QEC-like circuits, little to no advantage was found using gate-parallelization.
-QEC runtime is dominated by the time to collect mid-circuit measurements, and so improving the small chunks of non-measurement gates in between does not do much for the run time.
-However, it is included in STAB-Sim because there is no reason why a better suited workload could not take advantage of $sim2d$ to speed up simulation time if the user is interested in very many Clifford gates.
-3.3. Pauli Grouping Using STAB-Sim
-Pauli grouping is a well used and understood method in quantum chemistry and optimization.
-Because of this usefulness, and applications to algorithms 3 and 4 in our T-Gate transpiler, we quickly discuss functions for grouping in the context of binarized stabilizers.
-Using $x$ and $z$ as the bits of the $X$ and $Z$ components of the stabilizers (Pauli strings), $(x^{i}_{1}\land z^{i}_{2})\oplus(x^{i}_{2}\land z^{i}_{1})=0$ if they commute at each qubit $i$.
-Running an xor product can extend the commutation check from qubit-wise to group-wise over the full Pauli string:
-$\bigoplus_{i=1}^{n}((x^{i}_{1}\land z^{i}_{2})\oplus(x^{i}_{2}\land z^{i}_{1}))=0$
-if $P_{1}$ and $P_{2}$ commute.
-If the binary vectors can be packed into larger data types (e.g.
-32, 64-bit words), they can be quickly processed with vectorized bitwise instructions.
-For group-wise commutation, the xor product reduction can be replaced in favor of a population count on the vectorized results if its an available intrinsic.
-For example, if
-$\mathrm{popcount}\left(\left\{i\in[n]\;\middle|\;[P_{1}^{(i)},P_{2}^{(i)}]\right\}\right)\bmod 2=0,$
-then $P_{1}$ and $P_{2}$ group-commute.
-To evaluate the commutation of a Pauli string $P$ against an entire stabilizer tableau using matrix operations, the following matrix expression for all stabilizers $S$ in the tableau is used:
-$(x_{P}\cdot Z_{S})+(z_{P}\cdot X_{S})\mod 2=0$
-if $S$ commutes with the full tableau.
-Here, $x_{P}$ and $z_{P}$ are the binary row vectors representing the $X$ and $Z$ components of the Pauli string $P$.
-$X_{S}$ and $Z_{S}$ are the binary matrices corresponding to the $X$ and $Z$ components of all stabilizers $S$ in the tableau.
-The final sum matrix is an $n\times n$ matrix of the qubit-wise commutation results from $[P,S_{i}]$ where $i$ is each row.
-3.4. Clifford Push Through Transpiler
-Transpiling a Clifford+T universal quantum circuit into Pauli measurements with a separate T-layers has great potential to make efficient use of quantum machinery by abstracting away Clifford operations, and minimizing expensive magic state operations.
-Commuting T-layers can be optimized after separation by removing redundant eighth rotations and pushing them to the Clifford layer as quarter rotation gates.
-3.4.1. Measurement Tableau and T-Tableau Construction
-Algorithm 2 Tableau Construction
-1:Initialize an empty tableau $T_{tab}$ to store $pi/4$ rotation gates ($T$ and $T^{\dagger}$), and an $n\times n$ identity tableau $M_{tab}$.
-2:for all gates $G$ at qubit $q$ starting from the end do
-3:     if $G=T_{q}$ then
-4:         Append an $n$-qubit $Z_{q}$-stabilizer to $T_{tab}$ and set $r\leftarrow 0$
-5:     else if $G=T^{\dagger}_{q}$ then
-6:         Append an $n$-qubit $Z_{q}$-stabilizer to $T_{tab}$ and set $r\leftarrow 1$
-7:     else if $G$ is Clifford then
-8:         Apply $G$ to $M_{tab}$ and $T_{tab}$ using CHP rules
-9:     end if
-10:end for
-With the target Clifford+T circuit arranged in time order, gates are extracted from the end and added to one of two sets of tableaus.
-The first tableau is an n-qubit identity tableau of only stabilizers, used for tracking Clifford operations absorbed into the measurement.
-The second is an empty n-qubit tableau used for tracking the commutation rules of T gates as they are pushed through the circuit.
-If the gate extracted is a T gate, it is appended to the end of the T tableau as an n-qubit Pauli string with a Z stabilizer at the qubit index of the T gate.
-If the gate extracted is a Clifford operation, it is applied to the measurement (M) tableau by CHP gate operations.
-It is also applied to the T tableau by the same rules, but only acts on the T gates which have been appended up to that point.
-For example, if from the end of the original circuit, no T gates have been extracted yet, then the Clifford gate will have no affect on the T tableau as there are no interactions with the T gates that come before it in time order.
-Figure 4.
-Gate-by-gate extraction from the reversed circuit to construct the Measurement and T Tableaus.
-The Measurement tableau is a standard CHP tableau.
-The T-tableau treats T-gates as Z-stabilizers at the T-gate target qubit.
-Clifford gates are applied to any stabilizer that has been appended at the point where the gate occurs in reverse-time order.
-3.4.2. T-Layer Separation
-Algorithm 3 T-Separation
-1:Initialize tableau list $\mathcal{P}\leftarrow[\ ]$
-2:Create tableau $P_{0}$ and append the last stabilizer in $T_{\text{tab}}$
-3:Append $P_{0}$ to $\mathcal{P}$
-4:for all stabilizers $S$ from second-to-last to first in $T_{\text{tab}}$ do
-5:     $appended\leftarrow\textbf{false}$
-6:     for all $P\in\mathcal{P}$ starting from $P_{0}$ do
-7:         if $[S,P]=0$ then
-8:              Append $S$ to $P$
-9:              $appended\leftarrow\textbf{true}$
-10:              break
-11:         end if
-12:     end for
-13:     if not $appended$ then
-14:         Create new tableau $P_{\lvert\mathcal{P}\rvert}$ and append $S$
-15:         Append $P_{\lvert\mathcal{P}\rvert}$ to $\mathcal{P}$
-16:     end if
-17:end for
-Once the full circuit has been parsed and M and T tableaus have been created, the T tableau is split into commuting groups as follows.
-Starting from the end of the T tableau (the last appended stabilizer, i.e.
-the first T gate in time order) a new empty tableau that we will call $P_{0}$ is made and the stabilizer is appended.
-Each subsequent stabilizer from the end of the T tableau is checked against the new P tableau using the commutation method discussed in  3.3.
-If the subsequent stabilizer commutes with each Pauli string in the first P tableau, it is appended to that tableau.
-If there is an anti-commutation between the stabilizer being checked and a stabilizer within the tableau, a new tableau ($P_{1}$) is created.
-For the next stabilizer from the end of T, it is first checked against $P_{0}$.
-If it commutes with $P_{0}$ it is appended to $P_{0}$.
-If not, the stabilizer is checked against $P_{1}$ and appended to $P_{1}$ if it commutes.
-If it anti-commutes, the stabilizer starts a new tableau $P_{2}$.
-This process continues from the end of T for all stabilizers in T until T is empty.
-After this process is complete, there will be some number of self-commuting P tableaus (at least one, at most the number of T gates in the original circuit).
-Figure 5.
-T-Separation.
-The T-Gate tableau is separated into commuting groups by placing the first T-stabilizer at the end of the original circuit order.
-If the stabilizer does not commute, it cannot be grouped into the same T-layer, and must be appended to an earlier T-Layer or create its own.
-At this point, the basic measurement circuit and T layers have been created.
-The total number of T-stabilizers is counted to later determine when optimization is done.
-Clifford rotations are now extracted from the P tableaus with the goal of reducing the depth of the T layers.
-To do this, repeated stabilizers within each tableau are counted, starting from the last P tableau.
-The P tableaus are in reverse time order, meaning the first entry of $P_{0}$ contains the very last T gate in the original circuit.
-We want to push Clifford gates to the end, towards the measurement circuit, to absorb any leftover Clifford rotation.
-For every two occurrences of a stabilizer, a quarter rotation gate can be extracted.
-Likewise, for every four occurrences of a quarter rotation gate, a full rotation is made, and all eight T stabilizers can be removed without pushing any operations through to the measurement tableau.
-3.4.3. T-Layer Optimization
-Algorithm 4 T-Optimization
-Count initial total number of rows in all $P_{i}\in\mathcal{P}$, store as $N_{\text{prev}}$
-Initialize convergence flag $converged\leftarrow\textbf{false}$
-while not $converged$ do
-     for $i=\lvert\mathcal{P}\rvert-1$ down to 00 do
-         while there exist two identical stabilizers $S$ in $P_{i}$ do
-              Remove both copies of $S$ from $P_{i}$
-              Append one copy of $S$ to $P_{i+1}$
-              for all stabilizers $S_{T}$ in $P_{i+1}$ do
-                  if $[S_{T},S]\neq 0$ then
-                       Apply $\text{rowsum}(T,S)$
-                  end if
-              end for
-              Remove $S$ from $P_{i+1}$ and append to $P_{i+2}$
-         end while
-     end for
-     Count the total number of rows in all $P_{i}\in\mathcal{P}$, store as $N_{\text{new}}$
-     if $N_{\text{new}}=N_{\text{prev}}$ then
-         $converged\leftarrow\textbf{true}$
-     else
-         $N_{\text{prev}}\leftarrow N_{\text{new}}$
-     end if
-end while
-$\lvert\mathcal{P}\rvert$ is the number of $\pi/4$ rotation gate layers after minimization.
-The sum of all non-Identity Pauli operators in all $P_{i}$ stabilizers is the number of $\pi/4$ rotation gates.
-We start by rightward gate pushing from the last P tableau, i.e., the first T gates in original circuit order.
-Repeat stabilizers in the tableau are counted, removed from the tableau, and the corresponding quarter rotations are then checked For the second to last P tableau, repeat stabilizers are once again counted and removed, but now need to be commuted through the P tableau to the right before being added to the beginning of the measurement circuit.
-To do this, one copy of the repeated stabilizer is appended to the next P tableau and checked to see if it commutes with every stabilizer in the tableau.
-If it does, it is passed through to the next tableau.
-If not, a $rowsum$ 1 is called on the appended stabilizer with each row of the tableau that does not commute.
-$rowsum$ accumulates commutation the effect of an anti-commuting Pauli string throughout the stabilizer tableau by summing the contributions of each stabilizer Pauli and corresponding phase bit $r$ (Aaronson and Gottesman, 2004).
-The sum modulo 4 of the appended Pauli string respective non-commuting stabilizer becomes the new phase bit $r$ of that stabilizer.
-$rowsum$ is originally used in the CHP formalism to propagate the effect of measurement on the tableau.
-To satisfy the commutation rules for pushing a Clifford rotation gate through another arbitrary rotation (Litinski, 2019), a phase of $i$ is added to the $rowsum$ product.
-Figure 6. T Optimization. Every two repeating stabilizers are removed from the tableau and pushed left, following the $rowsum+i$ routine for each unique Pauli string.
-Once a pass of Algorithm 4 is complete, the total T-stabilizers remaining are counted.
-If no repeat stabilizers were extracted during optimization, the process ends.
-Otherwise, optimization is repeated to extract new repeat stabilizers created by $rowsum+i$, until the total number of T-stabilizers no longer changes.
-After optimization, the $P$ tableaus can be recombined to a larger $T$ tableau.
-The stabilizers of the $T$ tableau are the self-entangled layers of the $\pi/4$ rotation gates that remain.
-In Figure 2, every non-Identity Pauli operator in all the layers is counted as a T-rotation, for comparison with the number states needed for T-gates in the original circuit.
-4. Evaluation
-CPU simulator evaluations were all performed on an AMD EPYC 7502 compute node with a 3.35 GHz boost clock and 256 GB of RAM.
-GPU evaluations were performed on an NVIDIA H100 SXM node.
-These platforms were chosen for a stable and consistent test environment versus a laptop or desktop that may have power saving modes and background processes.
-Tests for STABSim-CPU and STABSim-GPU were compiled with GCC 10.3.0, and STABSim-GPU used CUDA 12.1.
-4.1. QEC Simulation
-Stabilizer simulation is made feasible in time and size by the relatively slow growth of the stabilizer tableau as qubit count increases compared to other simulation methods.
-However, measurement operations are $O(n^{2})$ worst-case in the original CHP formalism, one dimension due to every stabilizer potentially being checked, and another from every column participating in $rowsum$ when the row check conditions are met.
-Measurements greatly bottleneck the time for dynamic circuits, especially QEC codes where many repeated stabilizer measurements are the primary interest for evaluation.
-Stim improves on the deterministic measurement worst case, and, combined with 256-bit SIMD instructions, we find that Stim outperforms STAB-GPU in surface code up to distance 33 with 2179 qubits/tableau columns.
-Beyond distance 33, STAB-GPU scales better than existing stabilizer simulators due to the lower extra cost associated with adding a qubit to the simulation 7.
-The main constraint in STAB-GPU is the synchronization necessary for correctness, especially in measurement, which leads to a closer-to-linear increase in simulation time as a function of circuit size.
-STAB-CPU is implemented as a boolean tableau without bit-packing or other methods to speed up wide circuits.
-This allows STAB-CPU to be very lightweight and extensible.
-Startup time for STAB-CPU is less noticeable than STAB-GPU, and it has the edge in small simulations where the overhead for launching a CUDA kernel has a meaningful impact on simulation time.
-Figure 7. Full single-shot tableau simulation
-of planar surface codes as code distance increases. The structure follows the example given in 2.4.
-Gate-wise parallelization was also tested for a planar surface code.
-In that case, we found that even in very large tableaus, simulation time remained almost the exact same due to the time-dependence of measurement operations.
-Time improvement gained by parallelized qubit initialization and entanglement between many data and ancilla qubits was minimal compared to the impact of measurement.
-Figure 7 shows how different simulators scale in a less structured workload that still mimics QEC.
-In this scenario, half of the qubits have a Hadamard or phase gate applied at random.
-The initialized qubits in the first half are used as the control for a CX gate to each corresponding qubit in the second half.
-20% of the second half qubits are then measured, chosen at random.
-This cycle is layered $log_{2}(n)$ times for the sake of having a non-trivially deep circuit while keeping reasonable testing times, with new randomly chosen values for each layer.
-Figures 7 and  8 display similar scaling trends, with a tradeoff from Stim to STAB-GPU at $\approx 2000$ qubits and around the 10 second mark for simulation time.
-Qiskit-Aer method=’stabilizer’ stays closer to Stim in the random test, in part because Stim’s time complexity improvement in deterministic measurements is utilized for only about half of the measurements, versus nearly all measurements in surface code.
-STABSim-CPU has similar scaling concerns as Stim, but performs well on the compute node to a surprising extent.
-Stim should perform far fewer row operations per Clifford gate than STABSim-CPU due to the packed 256-bit AVX instructions it uses.
-One possibility is that frequent measurement operations over thousands of qubits require enough bit masking operations to outweigh the time saved on simple Clifford gates.
-One benefit the testing platform has over a standard laptop CPU is a large 128 MB L3 cache.
-Fewer cache misses could potentially lead to less gain for Stim from an efficient data layout, even in very large tableaus.
-Figure 8. Full single-shot tableau simulation
-of partially random circuit.
-Half of the qubits are initialized with Hadamard or phase gates at random, and control a corresponding qubit in the second half.
-$20\%$ of the second half qubits are measured.
-The process is repeated $log_{2}(n)$ times with new random gate applications.
-4.2. Pauli Grouping
-To demonstrate the utility of STABSim, we apply STABSim to Pauli grouping problems.
-We focus on grouping the Pauli strings from molecular Hamiltonians, which are widely used in quantum chemistry applications.
-The Hamiltonians were chosen for varying degrees of freedom and measurement gate sizes, while remaining practical to simulate with our state vector simulator.
-The detailed information about the molecular configuration settings and the Pauli group terms are shown in Table 2.
-To group the Pauli strings, the following steps were employed.
-First, Pauli terms are sorted up front by the weight of the Pauli term coefficients.
-Groups of commuting Pauli terms are then greedily constructed starting with the highest weight terms by either QWC or GC.
-This results in higher weight terms being grouped earlier so sampling resources can be prioritized to get a higher fidelity from more shots on larger terms.
-Table 2 counts the total number of grouped gates.
-To verify the correctness of the ground state energy result after grouping, all molecular Hamiltonians were evaluated with a statevector simulator we have developed.
-Our tableau-based simulator enables fast and scalable implementation of both strategies with xor product commutation checks.
-The time spent performing grouping and generating measurement settings was negligible compared to the overall simulation time, especially when considering the time for statevector-based energy verification.
-VQE in particular still faces challenges in exponential post-processing complexity (Scriva et al., 2024), but Pauli grouping is a useful and scalable pre-processing step that can be used with any Pauli-term based quantum algorithm to reduce resource overhead.
+A.3. Circuit Optimization Results
+Here we include extended evaluation data for Clifford+T benchmark circuits.
+Table 3 lists the detailed Clifford+T gate counts and relative speedup of STABSim versus our own Python reference, both using Algorithms 34, and 5.
+A more concise comparison to PyZX is provided in Figure 9.
+Circuit Python Time Total Gates T-Gates: Opt-T T-Reduction
+Adder N10 14.26x 189 88:16 5.5x
+Adder N4 34.61x 25 4:4 1.0x
+Adder N433 2.74x 6196 1536:768 2.0x
+Basis Test N4 51.88x 577 156:132 1.18x
+Basis Trotter N4 16.64x 9819 1860:1578 1.18x
+Bell N4 17.19x 386 70:68 1.03x
+Dnn N16 8.17x 30016 6656:5920 1.12x
+Dnn N2 46.46x 3414 760:672 1.13x
+Dnn N8 16.56x 15008 3328:2960 1.12x
+Fredkin N3 19.28x 22 4:4 1.0x
+Gcm H6 7.97x 17719 4852:4486 1.08x
+Inverseqft N4 13.29x 99 28:28 1.0x
+Ising N10 30.14x 3152 974:928 1.05x
+Ising N26 43.93x 1116 332:324 1.02x
+Knn N25 11.79x 816 204:192 1.06x
+Multiplier N15 3.32x 454 144:42 3.43x
+Multiplier N45 6.77x 4574 1512:396 3.82x
+Multiply N13 11.67x 92 24:14 1.71x
+Pea N5 13.69x 523 166:166 1.0x
+Qaoa N3 28.38x 53 8:8 1.0x
+Qaoa N6 28.65x 2928 672:544 1.24x
+Qec En N5 27.47x 25 1:1 1.0x
+Qf21 N15 18.42x 646 170:162 1.05x
+Qft N18 19.05x 2150 596:596 1.0x
+Qft N18 Iter1 29.37x 3290 1677:1333 1.26x
+Qft N18 Iter2 20.16x 27292 14383:11463 1.25x
+Qft N4 38.83x 310 92:92 1.0x
+Qpe N9 53.14x 334 90:86 1.05x
+Qram N20 32.09x 276 80:48 1.67x
+Sat N7 32.62x 213 40:24 1.67x
+Seca N11 29.27x 200 32:16 2.0x
+Simon N6 21.12x 56 8:8 1.0x
+Swap Test N115 10.99x 2866 736:680 1.08x
+Swap Test N25 13.76x 592 152:140 1.09x
+Teleportation N3 56.61x 8 1:1 1.0x
+Toffoli N3 22.36x 20 3:3 1.0x
+Variational N4 51.23x 363 100:92 1.09x
+Vqe N4 28.66x 639 188:186 1.01x
+Vqe Uccsd N4 15.41x 952 160:160 1.0x
+Vqe Uccsd N8 4.48x 33969 3816:3742 1.02x
+Wstate N27 26.17x 1704 416:416 1.0x
+Wstate N3 71.54x 67 14:14 1.0x
+Wstate N76 3.42x 2488 416:416 1.0x
 Table 3.
-Gates required for evaluation after grouping small molecular Hamiltonians.
-Initial basis gates, Qubit-Wise, and Group-Wise gates are shown, along with the % improvement from QWC to GWC in evaluation gate count.
-Molecule Orbitals Initial QWC GWC % Change
-H4 4 3428 717 275 61.65%
-LiH 3 756 449 316 29.62%
-LiH 4 3428 878 427 51.37%
-H6 18 16196 3844 1858 51.66%
-H6 22 16196 3850 1946 49.45%
-BeH2 6 16196 1681 749 55.44%
-BeH2 7 27124 3627 1580 56.44%
-4.3. PBC Transpilation
-Like Pauli grouping, the T-Transpiler application relies on fast commutation checks to be efficient.
-Figure 9 shows the time to perform Algorithm 3 to unpack an example circuit, and the time to perform optimization Algorithm 4, including the greedy commutation grouping described in Section 3.4.3.
-Similar to the Pauli grouping use-case, the purpose of the T-transpiler is to reduce quantum resources.
-Instead of reducing rotations for measurement settings, the quantity being optimized is the number of T-states in a static circuit along with trivializing the Clifford gates as rotations of the measurement basis.
-Figure 9. Time to process and optimize the circuit for select circuits using Python to track commutations as boolean values, versus STABSim.
-Circuit Python vs Clifford+T T- Optimized T-Ratio
-STAB Time Gates Gates Gates
-Adder N10 14.26x 189 88 16 2.0x
-Adder N4 34.61x 25 4 4 1.0x
-Adder N433 2.74x 6196 1536 768 2.0x
-Basis Test N4 51.88x 577 156 132 1.18x
-Basis Trotter N4 16.64x 9819 1860 1578 1.18x
-Bell N4 17.19x 386 70 68 1.03x
-Dnn N16 8.17x 30016 6656 5920 1.12x
-Dnn N2 46.46x 3414 760 672 1.13x
-Dnn N8 16.56x 15008 3328 2960 1.12x
-Fredkin N3 19.28x 22 4 4 1.0x
-Gcm H6 7.97x 17719 4852 4486 1.08x
-Inverseqft N4 13.29x 99 28 28 1.0x
-Ising N10 30.14x 3152 974 928 1.05x
-Ising N26 43.93x 1116 332 324 1.02x
-Knn N25 11.79x 816 204 192 1.06x
-Multiplier N15 3.32x 454 144 42 3.43x
-Multiplier N45 6.77x 4574 1512 396 3.82x
-Multiply N13 11.67x 92 24 14 1.71x
-Pea N5 13.69x 523 166 166 1.0x
-Qaoa N3 28.38x 53 8 8 1.0x
-Qaoa N6 28.65x 2928 672 544 1.24x
-Qec En N5 27.47x 25 1 1 1.0x
-Qf21 N15 18.42x 646 170 162 1.05x
-Qft N18 19.05x 2150 596 596 1.0x
-Qft N18 Iter1 29.37x 3290 1677 1333 1.26x
-Qft N18 Iter2 20.16x 27292 14383 11463 1.25x
-Qft N4 38.83x 310 92 92 1.0x
-Qpe N9 53.14x 334 90 86 1.05x
-Qram N20 32.09x 276 80 48 1.67x
-Sat N7 32.62x 213 40 24 1.67x
-Seca N11 29.27x 200 32 16 2.0x
-Simon N6 21.12x 56 8 8 1.0x
-Swap Test N115 10.99x 2866 736 680 1.08x
-Swap Test N25 13.76x 592 152 140 1.09x
-Teleportation N3 56.61x 8 1 1 1.0x
-Toffoli N3 22.36x 20 3 3 1.0x
-Variational N4 51.23x 363 100 92 1.09x
-Vqe N4 28.66x 639 188 186 1.01x
-Vqe Uccsd N4 15.41x 952 160 160 1.0x
-Vqe Uccsd N8 4.48x 33969 3816 3742 1.02x
-Wstate N27 26.17x 1704 416 416 1.0x
-Wstate N3 71.54x 67 14 14 1.0x
-Wstate N76 3.42x 2488 416 416 1.0x
-Table 3.
-Clifford+T transpilation speedup using STABSim.
-The total Clifford+T gates are counted for comparison after optimization.
-T-ratio shows the reduction of one-eighth rotations from the initial number of T-type gates.
-Most of the substantial time gain found with STAB-Sim is an improvement in the optimization step.
-Processing is primarily an issue of circuit size, and does not require commutation checks or other applications that can be sped up by bitwise methods.
-Table 3 shows how static Clifford+T circuits change under Clifford absorption into measurement and T-count optimization.
-The T-stabilizers remaining after optimization comprise the entire remaining circuit, apart from the now-merged and rotated measurement operations at the end.
-Each stabilizer in the T-tableau after the full processing is an entangled layer of $\pi/4$ rotational gates in the X, Y, or Z basis, given by the Pauli operator on each qubit.
-A phase bit of 1 in the stabilizer means the basis rotation in that layer is negative.
-Interestingly, circuits that showed the least potential for T-State reduction in Table 3 had the best improvement in checking the reduction from STAB-Sim versus our Python reference, as seen in Figure 9.
-The circuits with more trivial T-States tended to converge quickly and eliminate many T-layers, which reduced the size and frequency of loops in Algorithm 4, where commutation checks become a bottleneck.
-The difference in small circuits can also likely be attributed to the compiled nature of C++, especially the four qubit adder, where the only work is to add up the Clifford rotations for the end of circuit measurement settings, and check if there is any entanglement between the minimal four T-gates.
-In our testing, the QFT circuits compiled into Clifford+T circuits by applying the Solovay-Kitaev algorithm (Dawson and Nielsen, 2006) had the fewest trivial T-gates, meaning they were often barred from combining with other T-rotations.
-The compiled multiplier circuits tended to eliminate T-roations quickly, and mostly converge after the first round of optimization.
-Similar to the optimization performed in Table 2, the T-transpiler has application where simulation resources are limited or fewer magic states speed up circuit run times on a quantum device.
-If T-gates are sufficiently eliminated from the circuit, there is also a lower barrier towards Clifford-dominated circuits with few T-states being simulated by stabilizer simulation, as described by Bravyi and Gosset (Bravyi and Gosset, 2016).
-5. Related Work
-Many projects have been developed on the concept of simulating a quantum system from its stabilizers since CHP was proposed by Aaronson and Gottesman (Aaronson and Gottesman, 2004).
-Cirq, Qiskit, and other projects all implement stabilizer simulators as functions in their libraries (Developers, 2021; Qiskit Development Team, 2021).
-More recently, the Stim project made large gains in QEC sampling via Pauli frame simulation after the initial stabilizer simulation is complete (Gidney, 2021).
-Cuda-QEC leverages GPU for accelerated decoding of error syndromes (NVIDIA Corporation & Affiliates, [n.
-d.]).
-However, it does not provide a simulation method for evaluating error correcting codes.
-Qua-SARQ performs GPU accelerated simulation of Clifford gates on a stabilizer tableau, to replace the work of Stim in the equivalence checking applications (Osama et al., 2025).
-Crucially, Qua-SARQ only simulates Clifford gates, which, given the similarity to large binary matrix operations, are a portion of stabilizer simulation that end up being straightforward to parallelize.
-Qua-SARQ does not address measurement gates, which turn out to be the most costly portion of simulation and the crux of syndrome extraction in QEC.
-Pauli measurements and Pauli-Based Computation (PBC) are discussed significantly by Litinski in the context of surface code operations for fault-tolerance (Litinski, 2019).
-Peres and Galvão recently expanded on PBC compilation with a standalone Python compiler, and space/time tradeoff using auxiliary qubits (Peres and Galvão, 2023).
-To our knowledge, our work is the first to formulate PBC in the context of a stabilizer tableau, where highly efficient bitwise stabilizer operations can be exploited for compilation.
-TQEC
-The stabilizer tableau in this work also provides a natural platform to simulate the rotated Pauli measurements that are pushed through the circuit.
-Combined with our statevector simulator used in Section 4.2, or using a few T-gate method similar to what was proposed by Bravyi and Gosset (Bravyi and Gosset, 2016), the full compiled Clifford+T circuit can be evaluated by simulation.
-6. Conclusion
-Large stabilizer simulations become more and more of interest as quantum computing moves towards intermediate scale and eventual fault tolerance.
-Applications benefiting from large scale stabilizer-based computing can be found in higher-distance error correcting codes, quantum verification, limited Clifford+T circuits, heterogeneous codes, and more (Bravyi et al., 2024b; Fowler et al., 2012b; Bravyi and Gosset, 2016; Stein et al., 2024).
-STAB-Sim shows promise in evaluating growing QEC codes or other large circuits in reasonable amounts of time by scaling better than any existing stabilizer simulator.
-We furthermore show how the extended capabilities of STAB-Sim directly benefit other domains beyond direct simulation.
-Work in these areas has also inspired ideas for future directions.
-For example, while tracking Clifford and T push-through is relatively straightforward for static circuits using a limited amount of tableaus, it would be interesting to investigate how effectively a dynamic circuit with mid-circuit measurements could be transpiled, and if such a transpiler would be beneficial.
-Application of our fast Clifford simulatior to large Clifford+T circuits using Bravyi and Gosset’s algorithm would also be interesting to explore what universal quantum circuits could potentially be simulated in practical time.
-Acknowledgment
-This research was supported by the Quantum Algorithms and Architecture for Domain Science Initiative (QuAADS), under the Laboratory Directed Research and Development (LDRD) Program at Pacific Northwest National Laboratory (PNNL).
-M.
-Wang and A.
-Li were supported by the U.S.
-Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (QSC).
-This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.
-This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S.
-Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No.
-DE-AC02-05CH11231.
-The Pacific Northwest National Laboratory is operated by Battelle for the U.S.
-Department of Energy under Contract DE-AC05-76RL01830.
+Clifford+T transpilation speedup using STABSim vs the ZX-Calculus optimizer in PyZX to achieve the same reduction (Kissinger and Wetering, 2020).
+T:Opt shows the initial and optimized number of T-type gates, and T-reduction shows the reduction factor.